### v1

# Transformers learn through gradual rank increase

Abstract
We identify incremental learning dynamics in transformers, where the difference between trained and initial weights progressively increases in rank. We rigorously prove this occurs under the simplifying assumptions of diagonal weight matrices and small initialization. Our experiments support the theory and also show that phenomenon can occur in practice without the simplifying assumptions.
1 Introduction
The transformer architecture achieves state of the art performance in various domains, yet we still lack a solid theoretical understanding of its training dynamics [36, 11, 27, 10]. Nevertheless, the theoretical toolbox has matured over the last years and there are promising new approaches. One important line of work examines the role that initialization scale plays on the trajectory taken by gradient descent [20, 9, 17, 28, 21, 33, 24]. When the weights are initialized small, it has been shown for simple networks that an incremental learning behaviour occurs, where functions of increasing complexity are learned in stages. This regime is known to be richer than the large-initialization regime [^1] In the large-initialization regime, deep learning behaves as a kernel method [20, 9]. Various separations with kernels are known for smaller initialization: e.g., [16, 1, 31]., but the incremental learning dynamics are difficult to analyze, and are so far understood only for extremely simple architectures. Can we apply this analysis to transformers? Namely:
Are there incremental learning dynamics when training a transformer architecture?
An obstacle is that past work on incremental learning has mainly studied linear networks [7, 4, 30, 26, 37, 21, 18], with one paper studying nonlinear 2-layer fully-connected networks [8]. In contrast, transformers have nonlinear attention heads that do not fall under previous analyses: given ${\bm{X}}\in\mathbb{R}^{n\times d}$, an attention head computes
$\mathrm{attention}({\bm{X}};{\bm{W}}_{K},{\bm{W}}_{Q},{\bm{W}}_{V},{\bm{W}}_{O})=\mathrm{smax}({\bm{X}}{\bm{W}}_{K}{\bm{W}}_{Q}^{\top}{\bm{X}}^{\top}){\bm{X}}{\bm{W}}_{V}{\bm{W}}_{O}^{\top}$ (1)
where ${\bm{W}}_{K},{\bm{W}}_{Q},{\bm{W}}_{V},{\bm{W}}_{O}\in\mathbb{R}^{d\times d^{\prime}}$ are trainable matrices, and the softmax is applied row-wise. A transformer is even more complex, since it is formed by stacking alternating layers of attention heads and feedforward networks, along with residual connections.
Main finding
Our main finding is that transformers exhibit incremental learning dynamics, where the difference between the trained and initial weights incrementally increases in rank. Our results have a theoretical component and an experimental component.
Theoretical contributions
For our theory, we study a simplification of the transformer architecture, where the attention head weights are diagonal matrices: i.e., in each attention head we have ${\bm{W}}_{K}=\mathrm{diag}({\bm{w}}_{K})$, where ${\bm{w}}_{K}\in\mathbb{R}^{d}$ are trainable weights, and similarly for ${\bm{W}}_{Q},{\bm{W}}_{V}$ and ${\bm{W}}_{O}$. We rigorously establish the training dynamics of this architecture under gradient flow when the initialization is small. We prove that dynamics occur in discrete stages: (1) during most of each stage, the loss plateaus because the weights remain close to a saddle point, and (2) at the end, the saddle point is quickly escaped and the rank of the weights increases by at most one.
This theoretical result on transformers follows from a general theorem characterizing the learning dynamics of networks $f_{\mathsf{NN}}$ that depend on the product of parameters ${\bm{u}},{\bm{v}}\in\mathbb{R}^{p}$ as
$f_{\mathsf{NN}}({\bm{x}};{\bm{u}},{\bm{v}})=h({\bm{x}};{\bm{u}}\odot{\bm{v}})\,,$ (2)
where ${\bm{x}}$ is the input, $\odot$ denotes the elementwise product, and $h$ is a smooth function.
Theorem 1.1 (Informal statement of incremental learning dynamics).
Let $f_{\mathsf{NN}}$ be a network of the form (2), and suppose that the weights are initialized very small: i.e., the entries of ${\bm{u}},{\bm{v}}$ are initialized on the order $\Theta(\alpha)$ for some small $\alpha>0$. Then the dynamics of gradient flow training effectively proceeds in discrete stages, each one lasting time $\Theta(\log(1/\alpha))$. In each stage, the number of nonnegligible entries of ${\bm{u}}\odot{\bm{v}}$ increases by at most one.
A transformer with diagonal weight matrices falls under this result when we only train the attention head weights. For example, if the transformer has one attention head, then we can take ${\bm{u}}=[{\bm{w}}_{K},{\bm{w}}_{V}]\in\mathbb{R}^{2d}$ and ${\bm{v}}=[{\bm{w}}_{Q},{\bm{w}}_{O}]\in\mathbb{R}^{2d}$ to be concatenations of the diagonal entries of the weights of the head; see Example 3.2 for more details and the extension to transformers with many heads. Then, using Theorem 1.1, we see that in each stage either ${\bm{W}}_{K}{\bm{W}}_{Q}^{\top}=\mathrm{diag}({\bm{w}}_{K})\mathrm{diag}({\bm{w}}_{Q})$ or ${\bm{W}}_{V}{\bm{W}}_{O}^{\top}=\mathrm{diag}({\bm{w}}_{V})\mathrm{diag}({\bm{w}}_{O})$ increases in effective rank by at most one.222 We also remark that Theorem 1.1 is interesting in its own right and may have other applications beyond transformers. In fact, it qualitatively recovers the incremental dynamics result of [7] when specialized to linear diagonal networks, i.e., when $f_{\mathsf{NN}}({\bm{x}};{\bm{u}},{\bm{v}})=\sum_{i=1}^{p}u_{i}v_{i}x_{i}$. Furthermore, it addresses an open question of [7] for proving incremental learning dynamics without assuming ${\bm{u}}={\bm{v}}$ at initialization.
Experimental contributions
In our experiments, we first validate our theoretical results, which require the simplifying assumptions of small initialization and diagonal weight matrices.
Then, we conduct experiments on vision transformers in settings closer to practice, without any of the assumptions required by our theoretical analysis. Perhaps surprisingly, we again observe incremental learning dynamics, even though the assumptions of the theory are not met. We observe that the difference between trained and initial weights has low rank, and also that the rank of this difference grows gradually during training; see Figure 1. The incremental nature of the dynamics is easier to see for ImageNet, since for CIFAR-10 the rank of the weight difference does not grow as much.
Figure 1: For an attention head in ViT trained on (a) CIFAR-10, and (b) ImageNet, we plot the normalized spectra of ${\bm{W}}_{K}{\bm{W}}_{Q}^{\top}$ at initialization (in red), and of the learned perturbations to ${\bm{W}}_{K}{\bm{W}}_{Q}^{\top}$ at different epochs (in green).
1.1 Related work
Relation to LoRA
We note an intriguing connection to the LoRA algorithm, where a pretrained base model is cheaply fine-tuned by training a low-rank perturbation of the weights [25, 5, 19]. The method is surprisingly powerful, and recently LoRA has been fundamental to allowing the open-source community to inexpensively fine-tune language models [32, 34]. On the other hand, in our work we observe that the trained weights are a low-rank perturbation of the initial weights due to the training dynamics, without having to apply an explicit rank constraint as in LoRA. This raises an exciting open question for future work: can we explain and improve algorithms like LoRA by better understanding and quantifying the incremental dynamics of large transformers?
Low-rank bias in nonlinear models
For 2-layer networks, it is known that low-rank bias in the weights emerges if the target function depends on a low-dimensional subspace of the input [1, 2, 13, 6, 29]. The results of [1, 2] are especially relevant, since they show that the rank of the weights increases in a sequential manner, determined by the “leap complexity” of the target function, which is reminiscent of our empirical observations on transformers. See also [14, 35] for more investigations of low-rank bias in 2-layer networks under different assumptions. For transformers, [38] report that empirically the trained weights (using default initialization) are not low-rank. This is consistent with our claim that the difference between initial and trained weights is low-rank, since the initial weights might not be low-rank.
Incremental learning dynamics
Several works prove incremental learning behaviour in deep linear networks when the initialization is small. [15] has shown that gradient descent dynamics on a 2-layer linear network with $L_{2}$ loss effectively solve a reduced-rank regression problem with gradually increasing rank. [18] prove a dynamical depth separation result, allowing for milder assumptions on initialization scale. [4, 30] show implicit bias towards low rank in deep matrix and tensor factorization. [26] show deep matrix factorization dynamics with small initialization are equivalent to a greedy low-rank learning (GLRL) algorithm. And [21] independently provides a similar description of the dynamics, but without requiring balanced initialization. Finally, [7, 22] overcome a technical hurdle from previous analyses by proving incremental learning for the entire training trajectory, rather than just the first stage. In contrast to our result, these prior works apply only to linear networks with certain convex losses, whereas our result applies to nonlinear networks. In order to make our extension to nonlinear networks possible, we must make stronger assumptions on the training trajectory, which we verify hold empirically. As far as we are aware, one other work on incremental learning handles nonlinear networks: [8] proves that a 2-layer network learns with a two-stage incremental dynamic; but that result needs the stylized assumption that all data points are orthogonal.
1.2 Paper organization
Sections 23, and 4 contain theoretical preliminaries, definitions of the models to which our theory applies, and our main theoretical result on incremental dynamics. Section 5 provides experiments which verify and extend the theory. Section 6 discusses limitations and future directions.
2 Preliminaries
We consider training a network $f_{\mathsf{NN}}(\cdot;{\bm{\theta}})$ parametrized by a vector of weights ${\bm{\theta}}$, to minimize a loss
$\mathcal{L}({\bm{\theta}})=\operatorname{\mathbb{E}}_{{\bm{x}},{\bm{y}}}[\ell({\bm{y}},f_{\mathsf{NN}}({\bm{x}};{\bm{\theta}}))]\,,$
where the expectation is over samples $({\bm{x}},{\bm{y}})\in\mathbb{R}^{d_{y}}\times\mathbb{R}^{d_{x}}$ from a training data distribution, and $\ell:\mathbb{R}^{d_{y}}\times\mathbb{R}^{d_{out}}\to\mathbb{R}$. Consider a solution ${\bm{\theta}}(t)$ to the gradient flow
${\bm{\theta}}(0)=\alpha{\bm{\theta}}_{0},~{}~{}~{}\frac{d{\bm{\theta}}}{dt}=-\nabla_{{\bm{\theta}}}\mathcal{L}({\bm{\theta}})$ (3)
where $\alpha>0$ is a parameter governing the initialization scale, that we will take very small. For our theory, we henceforth require the following mild regularity assumption on the loss and data.
Assumption 2.1 (Regularity of data distribution and loss).
The function $\ell({\bm{y}},{\bm{\zeta}})$ is continuously twice-differentiable in the arguments $[{\bm{y}},{\bm{\zeta}}]\in\mathbb{R}^{d_{y}+d_{out}}$. There exists $C>0$ such that almost surely the data is bounded by $\|{\bm{x}}\|,\|{\bm{y}}\|\leq C$.
The assumption on $\ell$ is satisfied in typical cases such as the square and the cross-entropy losses. The data boundedness is often satisfied in practice (e.g., if the data is normalized).
3 Neural networks with diagonal weights
Our theory analyzes the training dynamics of networks that depend on products of diagonal weight matrices. We use $\odot$ to denote elementwise vector product.
Definition 3.1.
A network $f_{\mathsf{NN}}$ is smooth with diagonal weights ${\bm{\theta}}=({\bm{u}},{\bm{v}})\in\mathbb{R}^{2p}$ if it is of the form
$f_{\mathsf{NN}}({\bm{x}};{\bm{\theta}})=h({\bm{x}};{\bm{u}}\odot{\bm{v}})$
where $h:\mathbb{R}^{d_{x}}\times\mathbb{R}^{p}\to\mathbb{R}^{d_{out}}$ is continuously twice-differentiable in its arguments in $\mathbb{R}^{d_{x}+p}$.
The assumption on $h$ precludes the use of the ReLU function since it is not continuously-differentiable. Otherwise the assumption is fairly mild since any $h$ can be used to express an architecture of any depth as long as the nonlinearities are twice-differentiable, which includes for example GeLUs (as used in ViT). We describe how to express a transformer with diagonal weights.
Example 3.2 (Transformer with diagonal weights).
Consider a transformer with $L$ layers and $H$ attention heads on each layer. The transformer output at layer $\ell$ is ${\bm{Z}}_{\ell}\in\mathbb{R}^{n\times d}$, which is given by ${\bm{Z}}_{0}={\bm{X}}$ and inductively for $\ell>0$ by
  • (Attention layer) $\tilde{{\bm{Z}}}_{\ell}={\bm{Z}}_{\ell-1}+\sum_{i=1}^{H}\mathrm{attention}({\bm{Z}}_{\ell-1};{\bm{W}}_{K}^{\ell,i},{\bm{W}}_{Q}^{\ell,i},{\bm{W}}_{V}^{\ell,i},{\bm{W}}_{O}^{\ell,i})$
  • (Feedforward layer) ${\bm{Z}}_{\ell}=\tilde{{\bm{Z}}}_{\ell}+\sigma(\tilde{{\bm{Z}}}_{\ell}{\bm{W}}_{A}^{\ell})({\bm{W}}_{B}^{\ell})^{\top}$ ,
where ${\bm{W}}_{K}^{\ell,i},{\bm{W}}_{Q}^{\ell,i},{\bm{W}}_{V}^{\ell,i},{\bm{W}}_{O}^{\ell,i}\in\mathbb{R}^{d\times d^{\prime}}$ are attention parameters, and ${\bm{W}}_{A}^{\ell},{\bm{W}}_{B}^{\ell}\in\mathbb{R}^{d\times d^{\prime}}$ are the feedforward parameters, and $\sigma$ is a continuously twice-differentiable activation.
Suppose that the only trainable parameters are the attention parameters, and that these are diagonal matrices: i.e., ${\bm{W}}_{K}^{\ell,i}=\mathrm{diag}({\bm{w}}_{K}^{\ell,i})$ for some ${\bm{w}}_{K}^{\ell,i}\in\mathbb{R}^{d}$, and similarly for the other attention parameters. Because of the structure of the attention head (1), the final output ${\bm{Z}}_{L}$ only depends on the attention parameters through the elementwise products ${\bm{w}}_{K}^{\ell,i}\odot{\bm{w}}_{Q}^{\ell,i}$ and ${\bm{w}}_{V}^{\ell,i}\odot{\bm{w}}_{O}^{\ell,i}$. In other words, we can write
${\bm{Z}}_{L}=h({\bm{X}};{\bm{u}}\odot{\bm{v}})\,,$
for vectors ${\bm{u}}=[{\bm{w}}_{K}^{\ell,i},{\bm{w}}_{V}^{\ell,i}]_{(\ell,i)\in[L]\times[H]}\in\mathbb{R}^{2dHL}$ and ${\bm{v}}=[{\bm{w}}_{Q}^{\ell,i},{\bm{w}}_{O}^{\ell,i}]_{(\ell,i)\in[L]\times[H]}\in\mathbb{R}^{2dHL}$, and some smooth model $h$, which fits under Definition 3.1.
4 Incremental learning in networks with diagonal weights
Any model $f_{\mathsf{NN}}$ with diagonal weights as in Definition 3.1 evolves under the gradient flow (3) as
$\frac{d{\bm{u}}}{dt}$ $={\bm{v}}\odot{\bm{g}}({\bm{\theta}}),\quad\frac{d{\bm{v}}}{dt}={\bm{u}}\odot{\bm{g}}({\bm{\theta}})\quad\mbox{ where }$ (4)
${\bm{g}}({\bm{\theta}})$ $=-\operatorname{\mathbb{E}}_{{\bm{x}},y}[D\ell(y,h({\bm{x}};{\bm{u}}\odot{\bm{v}}))^{\top}Dh({\bm{x}};{\bm{u}}\odot{\bm{v}})^{\top}]\,.$
Here $D\ell({\bm{y}},\cdot)\in\mathbb{R}^{1\times d_{out}}$ is the derivative of $\ell$ in the second argument and $Dh({\bm{x}},\cdot)\in\mathbb{R}^{d_{out}\times p}$ is the derivative of $h$ in the second argument. We show that if initialization scale of ${\bm{\theta}}=({\bm{u}},{\bm{v}})$ is small, then learning proceeds in incremental stages, as given in Algorithm 1, where in each stage the effective sparsity of ${\bm{u}}$ and ${\bm{v}}$ increases by at most one.
4.1 Intuition for incremental learning dynamics
We develop an informal intuition for the result. First, we observe a conservation law that simplifies the dynamics. It can be viewed as the balancedness property for networks with linear activations [3, 12], specialized to the case of diagonal layers.
Lemma 4.1 (Conservation law).
For any $i\in[p]$ and any time $t$, we have
$u_{i}^{2}(t)-v_{i}^{2}(t)=u_{i}^{2}(0)-v_{i}^{2}(0)\,.$ (5)
Proof.
This follows from $\frac{d}{dt}(u_{i}^{2}-v_{i}^{2})=u_{i}v_{i}g_{i}({\bm{\theta}})-u_{i}v_{i}g_{i}({\bm{\theta}})=0$. ∎
This reduces the degrees of freedom and means that we need only keep track of $p$ parameters in total. Specifically, if we define $w_{i}(t):=u_{i}(t)+v_{i}(t)$, then the vector ${\bm{w}}={\bm{u}}+{\bm{v}}$ evolves by
$\frac{d{\bm{w}}}{dt}={\bm{w}}\odot{\bm{g}}({\bm{\theta}})\,.$ (6)
Using the conservation law (5), one can compute ${\bm{\theta}}(t)$ from ${\bm{w}}(t)$, so it remains to analyze the dynamics of ${\bm{w}}(t)$.
4.1.1 Stage 1 of dynamics
Stage 1A of dynamics: loss plateau for time $\Theta(\log(1/\alpha))$
At very early times $t$, we have ${\bm{\theta}}(t)\approx{\bm{0}}$ because the weights are initialized to be very small. Thus, we can approximate ${\bm{g}}({\bm{\theta}}(t))\approx{\bm{g}}({\bm{0}})$ and so we can solve for the evolution of ${\bm{w}}$:
${\bm{w}}(t)\approx{\bm{w}}(0)\odot e^{{\bm{g}}({\bm{0}})t}.$
This approximation is valid until one of the entries of ${\bm{\theta}}(t)$ reaches constant size, which one can show happens around time $t\approx T_{1}\cdot\log(1/\alpha)$ for
$T_{1}=\min_{i\in[p]}1/|g_{i}({\bm{0}})|\,.$
Until this time, the weights ${\bm{\theta}}(t)$ are small, the network remains close to its initialization, and so we observe a loss plateau.
Stage 1B of dynamics: nonlinear dynamics for time $O(1)$
Subsequently, we observe a rapid decrease of the loss and nonlinear dynamics during a $O(1)$-order time-scale. Indeed, suppose that the dynamics are “non-degenerate” in the sense that there is a unique coordinate $i_{0}$ such that $1/|g_{i_{0}}({\bm{0}})|=T_{1}$. Under this assumption, in stage 1A, the weights only grow significantly at coordinate $i_{0}$. So one can show that for any small $\epsilon>0$, there is a time $\underline{t}_{1}(\epsilon)\approx T_{1}\cdot\log(1/\alpha)$ such that $u_{i_{0}}(\underline{t}_{1})\approx\epsilon$, $v_{i_{0}}(\underline{t}_{1})\approx s\epsilon$ for some sign $s\in\{+1,-1\}$, and $|u_{i}(\underline{t}_{1})|,|v_{i}(\underline{t}_{1})|=o_{\alpha}(1)$ for all $i\neq i_{0}$.333 Without loss of generality, we can ensure that at initialization ${\bm{u}}(0)$ and ${\bm{u}}(0)+{\bm{v}}(0)$ are nonnegative. This implies ${\bm{u}}(t)$ is nonnegative. The fact that $u_{i_{0}}$ and $v_{i_{0}}$ are roughly equal in magnitude but might differ in sign is due to the conservation law (5). See Appendix C.3 for details.
Because all coordinates except for $i_{0}$ are negligibly small after stage 1A, we may perform the following approximation of the dynamics. Zero out the weights at coordinates except for $i_{0}$, and consider the training dynamics starting at $\tilde{{\bm{\theta}}}=(\epsilon{\bm{e}}_{i_{0}},s\epsilon{\bm{e}}_{i_{0}})$. After some constant time, independent of $\alpha$, these dynamics should approach a stationary point. Furthermore, all coordinates of ${\bm{u}}$ and ${\bm{v}}$ will remain zero except for the $i_{0}$ coordinate, so the sparsity of the weights will be preserved. In other words, we should expect there to be a time $\bar{t}_{1}=\underline{t}_{1}+O(1)\approx T_{1}\cdot\log(1/\alpha)$ such that
${\bm{\theta}}(\bar{t}_{1})\approx(a{\bm{e}}_{i_{0}},sa{\bm{e}}_{i_{0}}):={\bm{\theta}}^{1}\,,$
for some $a\in\mathbb{R}_{>0}$, such that ${\bm{\theta}}^{1}$ is a stationary point of the loss. [^4] The entries of ${\bm{u}}$ and ${\bm{v}}$ are close in magnitude (but may differ in sign) because of the conservation law (5). This is a good approximation because $\bar{t}_{1}-\underline{t}_{1}=O(1)$ is a constant time-scale, so the weights at coordinates except for $i_{0}$ remain negligible between times $\underline{t}_{1}$ and $\bar{t}_{1}$. Overall, we have argued that the network approximately reaches stationary point that is 1-sparse, where only the weights at coordinate $i_{0}$ are nonzero.
4.1.2 Later stages
We can extend the argument to any number of stages $k$, where in each stage the weights remain close to constant for time $\Theta(\log(1/\alpha))$ and then rapidly change during time $O(1)$, with the sparsity of the weights increasing by at most one. In order to analyze multiple stages, we must also keep track of the magnitude of the weights on the logarithmic scale because these evolve nonnegligibly throughout training. Inductively on $k$, suppose that there is some $T_{k}\in\mathbb{R},{\bm{b}}^{k}\in\mathbb{R}^{p}$ and ${\bm{\theta}}^{k}\in\mathbb{R}^{2p}$ and a time $\bar{t}_{k}\approx T_{k}\cdot\log(1/\alpha)$ such that
$\log_{\alpha}({\bm{w}}(\bar{t}_{k}))\approx{\bm{b}}^{k}\mbox{ and}{\bm{\theta}}(\bar{t}_{k})\approx{\bm{\theta}}^{k},$
where ${\bm{\theta}}^{k}$ is a stationary point of the loss. We argue for the inductive step that there is $T_{k+1}\in\mathbb{R}$ such that during times $t\in(\bar{t}_{k},T_{k+1}\cdot\log(1/\alpha)-\Omega(1))$ the weights remain close to the stationary point from the previous phase, i.e., ${\bm{\theta}}(t)\approx{\bm{\theta}}^{k}$. And at a time $\bar{t}_{k+1}\approx T_{k+1}\cdot\log(1/\alpha)$ we have
$\log_{\alpha}({\bm{w}}(\bar{t}_{k+1}))\approx{\bm{b}}^{k+1}\mbox{and }{\bm{\theta}}(\bar{t}_{k+1})\approx{\bm{\theta}}^{k+1},$
where ${\bm{\theta}}^{k+1}$ and ${\bm{b}}^{k+1}$ are defined below, and ${\bm{\theta}}^{k+1}$ is a stationary point of the loss whose support has grown by at most one compared to ${\bm{\theta}}^{k}$. The pseudocode for the evolution of ${\bm{b}}^{k}$ and ${\bm{\theta}}^{k}$ along the stages is given in Algorithm 1, and more details are provided below.
Algorithm 1 Incremental learning in networks with diagonal weights
1:  ${\bm{b}}^{0}$, ${\bm{\theta}}^{0}\leftarrow{\bm{0}}\in\mathbb{R}^{p}$, $T_{0}\leftarrow 0$
2:  for stage number $k=0,1,2,\ldots$ do
3:      # (A) Pick new coordinate $i_{k}\in[p]$ to activate.
4:     For each $i$, define time $\Delta_{k}(i)$ until active using (8).
5:     Pick winning coordinate $i_{k}$ using (9)
6:     Calculate time $T_{k+1}$ using (9) and break if $\infty$
7:     Update logarithmic weight approximation ${\bm{b}}^{k+1}$ using (10)
8:      # (B) Train activated coordinates to stationarity.
9:     ${\bm{\theta}}^{k+1}\leftarrow$ limiting dynamics point from (11)
10:  end for
Stage $(k+1)$A, loss plateau for time $\Theta(\log(1/\alpha))$
At the beginning of stage $k+1$, the weights are close to the stationary point ${\bm{\theta}}^{k}$, and so, similarly to stage 1A, linear dynamics are valid.
${\bm{w}}(t)\approx{\bm{w}}(\bar{t}_{k})\odot e^{{\bm{g}}({\bm{\theta}}^{k})(t-\bar{t}_{k})}\,.$ (7)
Using the conservation law (5), we derive a “time until active” for each coordinate $i\in[p]$, which corresponds to the time for the weight at that coordinate to grow from negligible to nonnegligible magnitude:
$\Delta_{k}(i)=\begin{cases}(b_{i}^{k}-1+\operatorname{sgn}(g_{i}({\bm{\theta}}^{k})))/g_{i}({\bm{\theta}}^{k}),&\mbox{ if }g_{i}({\bm{\theta}}^{k})\neq 0\\ \infty,&\mbox{ if }g_{i}({\bm{\theta}}^{k})=0\end{cases}$ (8)
The approximation (7) therefore breaks down at a time $t\approx T_{k+1}\cdot\log(1/\alpha)$, where
$T_{k+1}=T_{k}+\Delta_{k}(i_{k}),\quad i_{k}=\arg\min_{i\in[p]}\Delta_{k}(i)\,,$ (9)
which corresponds to the first time at the weights at a coordinate grow from negligible to nonnegligible magnitude. And at times $t\approx T_{k+1}\cdot\log(1/\alpha)$, on the logarithmic scale ${\bm{w}}$ is given by
$\log_{\alpha}({\bm{w}}(t))\approx{\bm{b}}^{k+1}:={\bm{b}}^{k}-{\bm{g}}({\bm{\theta}}^{k})\Delta_{k}(i_{k})\,,$ (10)
Stage $(k+1)$B of dynamics: nonlinear dynamics for time $O(1)$
Subsequently, the weights evolve nonlinearly during $O(1)$ time. To see this, if we make the non-degeneracy assumption that there is a unique coordinate $i_{k}$ such that $\Delta_{k}(i_{k})=\min_{i}\Delta_{k}(i)$, then this means that in stage $(k+1)$A, the only coordinate where weights grow from negligible to nonnegligible magnitude is $i_{k}$. Roughly speaking, for any $\epsilon>0$, there is a time $\underline{t}_{k+1}(\epsilon)\approx T_{k+1}\cdot\log(1/\alpha)$ such that
${\bm{\theta}}(\underline{t}_{k+1})\approx{\bm{\theta}}^{k}+(\epsilon{\bm{e}}_{i_{k}},\operatorname{sgn}(g_{i}({\bm{\theta}}^{k}))\epsilon{\bm{e}}_{i_{k}})\,,$
where the sign of the weights in coordinate $i_{k}$ comes from the conservation law (5). At this time, the weights are approximately the stationary point from stage $k$, plus a small perturbation. Consider the dynamics of ${\bm{\psi}}^{k}(t,\epsilon)\in\mathbb{R}^{2p}$ initialized at ${\bm{\psi}}^{k}(0,\epsilon)={\bm{\theta}}^{k}+(\epsilon{\bm{e}}_{i_{k}},\operatorname{sgn}(g_{i}({\bm{\theta}}^{k}))\epsilon{\bm{e}}_{i_{k}})$ and evolving according to the gradient flow $\frac{d{\bm{\psi}}^{k}(t,\epsilon)}{dt}=-\nabla_{{\bm{\theta}}}\mathcal{L}({\bm{\psi}}^{k})$. These dynamics may be highly nonlinear, so to control them let us assume that as we take $\epsilon$ to be small, they converge to a limiting point ${\bm{\theta}}^{k+1}$
$\lim_{\epsilon\to 0}\lim_{t\to\infty}{\bm{\psi}}^{k}(t,\epsilon)={\bm{\theta}}^{k+1}\,.$ (11)
Then we expect that at a time $\bar{t}_{k+1}=\underline{t}_{k+1}+O(1)\approx T_{k+1}\cdot\log(1/\alpha)$, we have ${\bm{\theta}}(\bar{t}_{k+1})\approx{\bm{\theta}}^{k+1}$. This concludes the inductive step.
4.2 Formal statement of incremental dynamics
We formally state our result. For ease of notation, we write ${\bm{\theta}}^{k}=({\bm{u}}^{k},{\bm{v}}^{k})$ and ${\bm{v}}^{k}={\bm{s}}^{k}\odot{\bm{u}}^{k}$ for some sign-flip vector ${\bm{s}}^{k}\in\{+1,-1\}^{k}$. This form of ${\bm{\theta}}^{k}$ can be guaranteed by the conservation law (5) of the dynamics; see Appendix C. We also denote $\mathrm{supp}({\bm{\theta}}^{k}):=\mathrm{supp}({\bm{u}}^{k})=\mathrm{supp}({\bm{v}}^{k})\subseteq[p]$.
We state our assumptions formally. First, we require that the dynamics be non-degenerate, in the sense that two coordinates do not become active at the same time. We also place a technical condition to handle the corner case when a coordinate leaves the support of active coordinates.
Assumption 4.2 (Nondegeneracy of dynamics in part (A)).
The initialization satisfies $u_{i}(0)\neq v_{i}(0)$ for all $i$. For stage $k$, either $T_{k+1}=\infty$ or there is a unique minimizer $i_{k}$ to $\min_{i}\Delta_{k}(i_{k})$ in (9). Finally, for all $i\in\mathrm{supp}({\bm{\theta}}^{k-1})\setminus\mathrm{supp}({\bm{\theta}}^{k})$ we have $g_{i}({\bm{\theta}}^{k})\neq 0$.
Next, we require that very small perturbations of the coordinates outside of $\mathrm{supp}({\bm{\theta}}^{k})$ do not change the dynamics. For this, it suffices that ${\bm{\theta}}^{k}$ be a strict local minimum.
Assumption 4.3 (Stationary points are strict local minima).
For stage $k$, there exist $\delta_{k}>0$ and $c_{k}>0$ such that for $\tilde{{\bm{u}}}\in B({\bm{u}}^{k},\delta)$ supported on $\mathrm{supp}({\bm{u}}^{k})$, we have
$\mathcal{L}(\tilde{{\bm{u}}},{\bm{s}}^{k}\odot\tilde{{\bm{u}}})\geq c_{k}\|{\bm{u}}^{k}-\tilde{{\bm{u}}}\|^{2}$
Finally, we require a robust version of the assumption (11), asking for convergence to a neighborhood of ${\bm{\theta}}^{k+1}$ even when the initialization is slightly noisy.
Assumption 4.4 (Noise-robustness of dynamics in part (B)).
For any stage $k$ with $T_{k+1}<\infty$ and any $\epsilon>0$, there are $\delta>0$ and $\tau:\mathbb{R}_{>0}\to\mathbb{R}$ such that the following holds. For any $\tilde{{\bm{u}}}\in B({\bm{u}}^{k},\delta)\cap\mathbb{R}_{\geq 0}^{p}$ supported on $\mathrm{supp}(\tilde{{\bm{u}}})\subseteq\mathrm{supp}({\bm{u}}^{k})\cup\{i_{k}\}$, there exists a unique solution ${\bm{\psi}}:[0,\infty)\to\mathbb{R}^{p}$ of the gradient flow $\frac{d{\bm{\psi}}}{dt}=-\nabla_{{\bm{\theta}}}\mathcal{L}({\bm{\psi}})$ initialized at ${\bm{\psi}}(0)=(\tilde{{\bm{u}}},{\bm{s}}^{k+1}\odot\tilde{{\bm{u}}})$, and at times $t\geq\tau(\tilde{\psi}_{i_{k}})$,
$\|{\bm{\psi}}(t)-{\bm{\theta}}^{k+1}\|<\epsilon\,.$
These assumptions are validated experimentally in Appendix A. Using them, we prove that incremental learning Algorithm 1 tracks the gradient flow dynamics if the initialization scale is small.
Theorem 4.5 (Incremental dynamics).
For any stage $k$ and time $t\in(T_{k},T_{k+1})$ the following holds under Assumptions 4.2, 4.3 and 4.4. There is $\alpha_{0}(t)>0$ such that for all $\alpha<\alpha_{0}$, there exists a unique solution ${\bm{\theta}}:[0,t\log(1/\alpha)]\to\mathbb{R}^{p}$ to the gradient flow (3) and
$\lim_{\alpha\to 0}{\bm{\theta}}(t\cdot\log(1/\alpha))\to{\bm{\theta}}^{k}\,\,,$
and at each stage the sparsity increases by at most one: $\mathrm{supp}({\bm{\theta}}^{k+1})\setminus\mathrm{supp}({\bm{\theta}}^{k})\subseteq\{i_{k}\}$.
Example 4.6 (Application: Incremental learning in diagonal transformer).
In Example 3.2, we showed that a diagonal transformer falls under Theorem 4.5. As a corollary, the gradient flow on a transformer with small initialization will learn in stages, where in each stage there will be at most one head $i\in[H]$ on one layer $\ell\in[L]$ such that either the rank of ${\bm{W}}_{K}^{\ell,i}({\bm{W}}_{Q}^{\ell,i})^{\top}=\mathrm{diag}({\bm{w}}_{K}^{\ell,i})\mathrm{diag}({\bm{w}}_{Q}^{\ell,i})$ or the rank of ${\bm{W}}_{V}^{\ell,i}({\bm{W}}_{O}^{\ell,i})^{\top}=\mathrm{diag}({\bm{w}}_{V}^{\ell,i})\mathrm{diag}({\bm{w}}_{O}^{\ell,i})$ increases by at most one.
5 Experimental results
We experimentally support our theoretical findings in a series of experiments: first on a toy model given by Equation 1, followed by experiments on a vision transformer on the CIFAR datasets. We defer additional experimental details and results to the appendix.
Toy models
We consider a toy model comprised of one self-attention layer with a single head as in (1), with either diagonal or full weight matrices. We initialize ${\bm{W}}_{K},{\bm{W}}_{Q},{\bm{W}}_{V},{\bm{W}}_{O}$ using Gaussian initialization with a small standard deviation, and train the model using GD on a regression task with 50-dimensional random Gaussian token inputs and targets from a teacher model. During training, we track the diagonal entries of ${\bm{W}}_{K}{\bm{W}}_{Q}^{\top}$ and ${\bm{W}}_{V}{\bm{W}}_{O}^{\top}$ in the diagonal case, and the singular values of ${\bm{W}}_{K}{\bm{W}}_{Q}^{\top}$ and ${\bm{W}}_{V}{\bm{W}}_{O}^{\top}$ in the full weights case. Our results are summarized in Figure 2. For the diagonal model, as predicted, diagonal components are learned incrementally, resulting in progressive increase in the rank; in Appendix A we run additional experiments to verify that the assumptions of Theorem 4.5 indeed hold. For the full-weights model, we also observe incremental learning with progressively-increasing rank, even though this setting falls beyond our theory.
Figure 2: A network containing a single self-attention layer with diagonal (a) - (c) and full (d) - (f) weight matrices, trained with gradient descent in the incremental learning regime. (a) The diagonal entries of ${\bm{W}}_{V}{\bm{W}}_{O}^{\top}$ and (d) the singular values of ${\bm{W}}_{V}{\bm{W}}_{O}^{\top}$ are learned incrementally. (b) The diagonal entries of ${\bm{W}}_{K}{\bm{W}}_{Q}^{\top}$ and (e) the singular values of ${\bm{W}}_{K}{\bm{W}}_{Q}^{\top}$ are learned incrementally. (c), (f) The loss curves show stagewise plateaus and sharp decreases.
Vision transformers
We next run experiments that go well beyond our toy model to test the extent to which incremental learning with a low-rank bias exists in popular models used in practice. We conduct experiments with vision transformers (ViT) [10] trained on the CIFAR-10/100 and ImageNet datasets. We use a ViT of depth 6, with 8 self-attention heads per layer (with layer normalization). We use an embedding and MLP dimension of $d_{\text{emb}}=512$, and a head dimension of $d_{h}=128$ (i.e ${\bm{W}}_{K},{\bm{W}}_{Q},{\bm{W}}_{V},{\bm{W}}_{O}\in\mathbb{R}^{d_{\text{emb}}\times d_{h}}$). We train the transformer using Adam on the CIFAR-10/100 and ImageNet classification tasks with cross-entropy loss. We train all layers (including the feedforward layers) while varying the initialization scale of all layers by multiplying their initial values by a scale factor (we fix the scale of the initial token mapper). To illustrate the effect of training on weights with a non-vanishing initialization scale, we plot the spectrum of the difference $\Delta{\bm{W}}_{K}{\bm{W}}_{Q}^{\top}$ and $\Delta{\bm{W}}_{V}{\bm{W}}_{O}^{\top}$ between the weights post-training, and their initial values. Figure 3 shows the evolution of the principal components of $\Delta{\bm{W}}_{K}{\bm{W}}_{Q}^{\top}$ and $\Delta{\bm{W}}_{V}{\bm{W}}_{O}^{\top}$ for a randomly-chosen self-attention head and layer throughout training, exhibiting incremental learning dynamics and a low-rank bias. Note that incremental learning and low-rank bias are increasingly evident with smaller initialization scales, as further demonstrated in Figure 4. Finally, we plot the spectrum of $\Delta{\bm{W}}_{K}{\bm{W}}_{Q}^{\top}$ against that of its initialized state in Figure 5 for different self-attention heads, illustrating that the weight perturbation learned during the training process is extremely low-rank when compared to the initial spectrum. All figures in this section are given for models trained on CIFAR-10. In the appendix we conduct further experiments on CIFAR-100 and ImageNet, as well as different model sizes for completeness, and these show similar trends. Further experimental details and results are provided in the appendix.
Figure 3: Training a vision transformer on CIFAR-10 using Adam, while varying the initialization scale (unit scale indicates default initialization). Plotted are the evolution of the eigenvalues of $\Delta{\bm{W}}_{K}{\bm{W}}_{Q}^{\top}$ (a) - (c) and $\Delta{\bm{W}}_{V}{\bm{W}}_{O}^{\top}$ (d) - (f) in a random self-attention head in the second layer throughout training. Incremental learning dynamics and a low-rank bias are evident for all scales, albeit more pronounced at smaller initialization scales.
Figure 4: Stable rank of $\Delta{\bm{W}}_{K}{\bm{W}}_{Q}^{\top}$ per initialization scale (Unit scale refers to the default initialization) in different self-attention heads post-training, at layers 1, 3, 5. At each layer, the stable rank mean and standard deviation are computed across 8 heads per layer, for each initialization scale. All models were trained on CIFAR-10 using the Adam optimizer. Smaller initialization scales lead to lower-rank attention heads.
Figure 5: Spectrum of the weight perturbation $\Delta{\bm{W}}_{K}{\bm{W}}_{Q}^{\top}$ vs. initialization in a vision transformer trained on CIFAR-10, using Adam and default initialization scale, in random self-attention heads in different layers. The learned perturbation exhibits extreme low-rank bias post-training even in default initialization scales.
6 Discussion
We have identified incremental learning dynamics in transformers, proved them rigorously in a simplified setting, and shown them experimentally in networks trained with practical hyperparameters.
Limitations
There are clear limitations to our theory: the diagonal weights and small initialization assumptions. More subtly, the theory does not apply to losses with exponential-like tails because the weights may not converge to a finite value and so Assumption 4.3 is not met (this could possibly be addressed by adding regularization). Also, the architecture must be smooth, which precludes ReLUs – but allows for smoothed ReLUs such as the GeLUs used in ViT [10]. Finally, the theory is for training with gradient flow, while other optimizers such as Adam are used in practice instead [23]. Nevertheless, our experiments on ViTs indicate that the incremental learning dynamics occur even when training with Adam.
Future directions
A promising direction of future research is to examine the connection between our results on incremental dynamics and the LoRA method [19], with the goal of explaining and improving on this algorithm; see also the discussion in Section 1.1. Another interesting avenue is to develop a theoretical understanding of the implicit bias in function space of transformers whose weights are a low-rank perturbation of randomly initialized weights.
Acknowledgments
We would like to thank Vimal Thilak for his help in setting up the infrastructure for conducting experiments.
Appendix A Experimental validation of the assumptions in Theorem 4.5
In Figures 6, 7, and 8, we plot the evolution of the losses, of the entries of ${\bm{W}}_{K}{\bm{W}}_{Q}^{\top}=\mathrm{diag}({\bm{w}}_{K})\mathrm{diag}({\bm{w}}_{Q})$, and of the entries of ${\bm{W}}_{V}{\bm{W}}_{O}^{\top}=\mathrm{diag}({\bm{w}}_{V})\mathrm{diag}({\bm{w}}_{O})$ in the toy task of training an attention head (1) with diagonal weights. The model is trained with SGD on the mean-squared error loss on 1000 random samples $({\bm{X}},{\bm{y}})$. Each random sample has ${\bm{X}}\in\mathbb{R}^{10\times 50}$, which a sequence of 10 tokens, each of dimension 50, which is distributed as isotropic Gaussians. The label ${\bm{y}}$ is given by a randomly-generated teacher model that is also an attention head (1) with diagonal weights. In Figures 6, 7, and 8, for $\alpha\in\{0.1,0.01,0.0001,10^{-8},10^{-16},10^{-32}\}$ we plot the evolution of the loss and of the weights when initialized at ${\bm{\theta}}(0)=\alpha{\bm{\theta}}_{0}$, for some random Gaussian ${\bm{\theta}}_{0}$. Qualitatively, as $\alpha\to 0$ we observe that the loss curve and the trajectories of the weights appear to converge to a limiting stagewise dynamics, where there are plateaus followed by movement on short time-scales, as predicted by the theory.
Validation of Assumption 4.2 (non-degeneracy of dynamics)
As $\alpha\to 0$, notice that the stages appear to separate and happen at distinct times. Furthermore, the extra technical condition on coordinates $i\in\mathrm{supp}({\bm{\theta}}^{k})\setminus\mathrm{supp}({\bm{\theta}}^{k-1})$ in Assumption 4.2 is satisfied since no coordinates ever leave the support of ${\bm{\theta}}^{k}$.
Validation of Assumption 4.3 (stationary points are strict local minima)
In Figure 9 we consider the $\alpha=10^{-32}$ trajectory, since this is closest to the dynamics in the $\alpha\to 0$ limit. We randomly select several epochs. Since the transitions between stages are a vanishing fraction of the total training time, each of these randomly-selected epochs is likely during a plateau, as we see in the figure. For each epoch perform the following experiment. For each nonnegligible coordinate of the weights (those where the weight is of magnitude greater than the threshold $\tau=10^{-5}$), we perturb the weights by adding noise of standard deviation $0.05$. We then run the training dynamics starting at this perturbed initialization for 1000 epochs. We observe that the training dynamics quickly converge to the original unperturbed initialization, indicating that the weights were close to a strict local minimum of the loss.
Validation of Assumption 4.4 (noise-robustness of dynamics)
In Figure 10 we perform the same experiment as in Figure 9, except that the epochs we select to perturb the weights are those where there is a newly-nonnegligible coordinate (less than $10^{-5}$ in magnitude in the previous epoch, and more than $10^{-5}$ in magnitude in this epoch). We find that the nonlinear dynamics are robust and tend to the limiting endpoint even under a random Gaussian perturbation of standard deviation $10^{-2}$ on each of the nonnegligible coordinates, supporting Assumption 4.4.
Figure 6: Evolution of loss versus rescaled time initializing at various scalings $\alpha$ in the toy task of learning an attention head with diagonal weights. The loss curves converge as $\alpha\to 0$ to a curve with loss plateaus and sharp decreases, as predicted by the theory.
Figure 7: Evolution of $\mathrm{diag}({\bm{w}}_{Q})\mathrm{diag}({\bm{w}}_{K})$ entries over rescaled time initializing at various scalings $\alpha$. Notice that as $\alpha\to 0$, the training trajectories tend to a limiting trajectory. Each line corresponds to a diagonal entry of $\mathrm{diag}({\bm{w}}_{Q})\mathrm{diag}({\bm{w}}_{K})$.
Figure 8: Evolution of $\mathrm{diag}({\bm{w}}_{V})\mathrm{diag}({\bm{w}}_{O})$ entries in the toy task of learning an attention head with diagonal weights. Each line corresponds to the evolution of an entry of $\mathrm{diag}({\bm{w}}_{V})\mathrm{diag}({\bm{w}}_{O})$ over rescaled time. Each plot corresponds to a different initialization magnitude $\alpha$. Notice that as $\alpha\to 0$, the training trajectories tend to a limiting trajectory.
Figure 9: Evolution of weights of toy attention model under perturbation, validating Assumption 4.3. At 5 different random times during training, we perturb the nonnegligible weight coordinates and continue to train with SGD. The evolution of each of the weights under the initial perturbation (solid line) is compared to the original evolution without perturbation (dashed line). Observe that the training dynamics quickly brings each weight back to the unperturbed weight trajectory, indicating that the weights are originally close to a strict local minimum.
Figure 10: Validating Assumption 4.4 with the same experiment as in Figure 9, except that the epochs for the perturbation chosen are those where there is a newly nonnegligible coordinate. Perturbed dynamics (solid lines) are again robust to perturbation and track the original dynamics (dashed lines).
Appendix B Further experiments on vision transformers
The practice of training transformer models often deviates substantially from the assumptions made in our theoretical analysis, and it is a priori unclear to what extent gradual rank increase behaviour and a low rank bias are manifested in setups more common in practical applications. To gauge the relevancy of our findings we conduct experiments on popular vision benchmarks, using algorithms and hyperparameters common in the literature. We use the stable rank given by $\frac{\|s\|_{F}^{2}}{\|s\|_{2}^{2}}$, where $s$ is the spectrum, as a smooth approximation of rank. We track the value of the stable rank for the different attention matrices throughout training. Although we do not expect our theoretical results to to hold precisely in practice, we find evidence of gradual increase in stable rank, leading to a low rank bias Figures 12, 14 and 16. In these experiments we use off the shelf vision transformers (ViT) [10] trained on popular vision benchmarks. For the CIFAR-10/100 datasets we use a VIT with 6 layers, patchsize of 4, 8 heads per self attention layer, an embedding and MLP dimension of 512, and a head dimension of 128. We train the model using the Adam optimizer for 500 epochs with a base learning rate of 1e-4, a cyclic learning rate decay with a linear warmup schedule for 15 epochs and a batchsize of 512. For ImageNet, we use the VIT-Base/16 from [10] trained with Adam for 360 epochs with a base learning rate of 3e-3, a cyclic learning rate decay with a linear warmup schedule for 15 epochs and a batchsize of 4096. We use no weight decay or dropout in our experiments. All models were initialized using the default initialization scale. Our results are summarized in Figures 11 and 12 for CIFAR-10, Figures 13 and 14 for CIFAR-100 and Figures 15 and 16 for ImageNet.
B.1 CIFAR-10
Figure 11: CIFAR-10: normalized spectrum at different stages of training. (a) - (c) Normalized spectrum of ${\bm{W}}_{K}{\bm{W}}_{Q}^{\top}$ at initialization and $\Delta{\bm{W}}_{K}{\bm{W}}_{Q}^{\top}$ during training for different attention heads at different layers. (d) - (e) equivalent figures for ${\bm{W}}_{V}{\bm{W}}_{O}^{\top}$.
Figure 12: CIFAR-10: Stable rank of $\Delta{\bm{W}}_{K}{\bm{W}}_{Q}^{\top}$ (blue) and $\Delta{\bm{W}}_{V}{\bm{W}}_{O}^{\top}$ (red) throughout training. Mean and standard deviation (shaded area) are computed across 8 heads per attention layer.
B.2 CIFAR-100
Figure 13: CIFAR-100: normalized spectrum at different stages of training. (a) - (c) Normalized spectrum of ${\bm{W}}_{K}{\bm{W}}_{Q}^{\top}$ at initialization and $\Delta{\bm{W}}_{K}{\bm{W}}_{Q}^{\top}$ during training for different attention heads at different layers. (d) - (e) equivalent figures for ${\bm{W}}_{V}{\bm{W}}_{O}^{\top}$.
Figure 14: CIFAR-100: Stable rank of $\Delta{\bm{W}}_{K}{\bm{W}}_{Q}^{\top}$ (blue) and $\Delta{\bm{W}}_{V}{\bm{W}}_{O}^{\top}$ (red) throughout training. Mean and standard deviation (shaded area) are computed across 8 heads per attention layer.
B.3 ImageNet
Figure 15: ImageNet: normalized spectrum at different stages of training. (a) - (c) Normalized spectrum of ${\bm{W}}_{K}{\bm{W}}_{Q}^{\top}$ at initialization and $\Delta{\bm{W}}_{K}{\bm{W}}_{Q}^{\top}$ during training for different attention heads at different layers. (d) - (e) equivalent figures for ${\bm{W}}_{V}{\bm{W}}_{O}^{\top}$.
Figure 16: ImageNet: Stable rank of $\Delta{\bm{W}}_{K}{\bm{W}}_{Q}^{\top}$ (blue) and $\Delta{\bm{W}}_{V}{\bm{W}}_{O}^{\top}$ (red) throughout training. Mean and standard deviation (shaded area) are computed across 12 heads per attention layer.
Appendix C Proof for dynamics of networks with diagonal parametrization (Theorem 4.5)
C.1 Assumptions
Recall we have defined ${\bm{\theta}}^{0},\ldots,{\bm{\theta}}^{k},\ldots\in\mathbb{R}^{2p}$ as the sequence of weights such that ${\bm{\theta}}^{0}={\bm{0}}$ and ${\bm{\theta}}^{k+1}$ is defined inductively as follows. Consider the dynamics of ${\bm{\psi}}^{k}(t,\epsilon)\in\mathbb{R}^{2p}$ initialized at ${\bm{\psi}}^{k}(0,\epsilon)={\bm{\theta}}^{k}+(\epsilon{\bm{e}}_{i_{k}},\operatorname{sgn}(g_{i}({\bm{\theta}}^{k}))\epsilon{\bm{e}}_{i_{k}})$ and evolving according to the gradient flow $\frac{d{\bm{\psi}}^{k}(t,\epsilon)}{dt}=-\nabla_{{\bm{\theta}}}\mathcal{L}({\bm{\psi}}^{k})$. We assume that there is a limiting point ${\bm{\theta}}^{k+1}$ of these dynamics as $\epsilon$ is taken small and the time is taken large:
$\lim_{\epsilon\to 0}\lim_{t\to\infty}{\bm{\psi}}^{k}(t,\epsilon)={\bm{\theta}}^{k+1}\,.$
Under the above assumption that this sequence ${\bm{\theta}}^{0},\ldots,{\bm{\theta}}^{k},\ldots$ is well-defined, we can derive a useful property of it for free. Namely, the conservation law (5) implies that ${\bm{u}}\odot{\bm{u}}-{\bm{v}}\odot{\bm{v}}$ is preserved. It follows that for each $k$ we have that ${\bm{\theta}}^{k}=({\bm{u}}^{k},{\bm{v}}^{k})$ satisfies $|{\bm{u}}^{k}|=|{\bm{v}}^{k}|$ entrywise. In other words, there is ${\bm{s}}^{k}\in\{+1,-1\}^{p}$ satisfying
${\bm{\theta}}^{k}$ $=({\bm{u}}^{k},{\bm{s}}^{k}\odot{\bm{u}}^{k})\in\mathbb{R}^{2p}\,.$
We also abuse notation and write $\mathrm{supp}({\bm{\theta}}^{k}):=\mathrm{supp}({\bm{u}}^{k})\subseteq[p]$, since the support of ${\bm{\theta}}^{k}$ on the first $p$ coordinates matches its support on the last $p$ coordinates.
Having fixed this notation, we now recall the main assumptions of the theorem.
Assumption C.1 (Nondegeneracy of dynamics in part (A); Assumption 4.2).
The initialization satisfies $u_{i}(0)\neq v_{i}(0)$ for all $i$. For stage $k$, either $T_{k+1}=\infty$ or there is a unique minimizer $i_{k}$ to $\min_{i}\Delta_{k}(i_{k})$ in (9). Finally, for all $i\in\mathrm{supp}({\bm{\theta}}^{k-1})\setminus\mathrm{supp}({\bm{\theta}}^{k})$ we have $g_{i}({\bm{\theta}}^{k})\neq 0$.
Assumption C.2 (Stationary points are strict local minima; Assumption 4.3).
For stage $k$, there exist $\delta_{k}>0$ and $c_{k}>0$ such that for $\tilde{{\bm{u}}}\in B({\bm{u}}^{k},\delta)$ supported on $\mathrm{supp}({\bm{u}}^{k})$, we have
$\mathcal{L}(\tilde{{\bm{u}}},{\bm{s}}^{k}\odot\tilde{{\bm{u}}})\geq c_{k}\|{\bm{u}}^{k}-\tilde{{\bm{u}}}\|^{2}\,.$
Assumption C.3 (Noise-robustness of dynamics in part (B); Assumption 4.4).
For stage $k$, either $T_{k+1}=\infty$ or the following holds. For any $\epsilon>0$, there are $\delta>0$ and $\tau:\mathbb{R}_{>0}\to\mathbb{R}$ such that the following holds. For any $\tilde{{\bm{u}}}\in B({\bm{u}}^{k},\delta)\cap\mathbb{R}_{\geq 0}^{p}$ supported on $\mathrm{supp}(\tilde{{\bm{u}}})\subseteq\mathrm{supp}({\bm{u}}^{k})\cup\{i_{k}\}$, there exists a unique solution ${\bm{\psi}}:[0,\infty)\to\mathbb{R}^{p}$ of the gradient flow $\frac{d{\bm{\psi}}}{dt}=-\nabla_{{\bm{\theta}}}\mathcal{L}({\bm{\psi}})$ initialized at ${\bm{\psi}}(0)=(\tilde{{\bm{u}}},{\bm{s}}^{k+1}\odot\tilde{{\bm{u}}})$, and at times $t\geq\tau(\tilde{u}_{i_{k}})$,
$\|{\bm{\psi}}(t)-{\bm{\theta}}^{k+1}\|<\epsilon\,.$
C.2 Rescaling time for notational convenience
For ease of notation, we rescale time
${\bm{u}}_{\alpha}(0)=\alpha{\bm{u}}(0),\quad{\bm{v}}_{\alpha}(0)=\alpha{\bm{v}}(0)$
$\frac{d{\bm{u}}_{\alpha}}{dt}=\log(1/\alpha){\bm{v}}_{\alpha}\odot{\bm{g}}({\bm{u}}_{\alpha},{\bm{v}}_{\alpha}),\quad\frac{d{\bm{v}}_{\alpha}}{dt}=\log(1/\alpha){\bm{u}}_{\alpha}\odot{\bm{g}}({\bm{u}}_{\alpha},{\bm{v}}_{\alpha}).$ (12)
We also define
${\bm{\theta}}_{\alpha}(t)$ $=({\bm{u}}_{\alpha}(t),{\bm{v}}_{\alpha}(t))\in\mathbb{R}^{2p}\,.$
Because of this time-rescaling, we equivalently state Theorem 4.5 as:
Theorem C.4 (Restatement of Theorem 4.5).
Let $K\in\mathbb{Z}_{\geq 0}$ be such that Assumptions 4.2  4.3 hold for all $k\leq K$ and Assumption 4.4 holds for all $k<K$. Then for any $k\leq K$ and time $t\in(T_{k},T_{k+1})$ the following holds. There is $\alpha_{0}(t)>0$ such that for all $\alpha<\alpha_{0}$, there exists a unique solution ${\bm{\theta}}_{\alpha}:[0,t]\to\mathbb{R}^{p}$ to the gradient flow (12) and
$\lim_{\alpha\to 0}{\bm{\theta}}_{\alpha}(t)\to{\bm{\theta}}^{k}\,,$
where at each stage $|\mathrm{supp}({\bm{u}}^{k})\setminus\mathrm{supp}({\bm{u}}^{k-1})|\leq 1$.
For shorthand, we also write
$S_{k}=\mathrm{supp}({\bm{u}}^{k})\mbox{ and }S_{k}^{c}=[p]\setminus\mathrm{supp}({\bm{u}}^{k})\,.$
C.3 Simplifying problem without loss of generality
For each coordinate $i\in[p]$ we have $|u_{\alpha,i}(0)|\neq|v_{\alpha,i}(0)|$ by the non-degeneracy Assumption 4.2. So we can assume $|u_{\alpha,i}(0)|>|v_{\alpha,i}(0)|$ without loss of generality. Furthermore, we can assume the entrywise inequality
${\bm{u}}_{\alpha}(0)>0$
by otherwise training weights $\tilde{{\bm{u}}}_{\alpha}(t),\tilde{{\bm{v}}}_{\alpha}(t)$ initialized at $\tilde{{\bm{u}}}_{\alpha}(0)=\operatorname{sgn}({\bm{u}}_{\alpha}(0)){\bm{u}}_{\alpha}(0)$ and $\tilde{{\bm{v}}}_{\alpha}(0)=\operatorname{sgn}({\bm{v}}_{\alpha}(0)){\bm{v}}_{\alpha}(0)$, as $\tilde{{\bm{u}}}_{\alpha}(t)\odot\tilde{{\bm{v}}}_{\alpha}(t)={\bm{u}}_{\alpha}(t)\odot{\bm{v}}_{\alpha}(t)$ at all times.
Since $u_{\alpha,i}^{2}(t)-v_{\alpha,i}^{2}(t)=u_{\alpha,i}^{2}(0)-v_{\alpha,i}^{2}(0)$ by the conservation law (5), it holds that $|u_{\alpha,i}(t)|>|v_{\alpha,i}(t)|$ throughout. So by continuity
${\bm{u}}_{\alpha}(t)>0$
throughout training.
C.4 Tracking the sum of the weights
We define
${\bm{w}}_{\alpha}(t)={\bm{u}}_{\alpha}(t)+{\bm{v}}_{\alpha}(t)\,.$
The reason for this definition is that during training we have
$\frac{d{\bm{w}}_{\alpha}}{dt}=\log(1/\alpha){\bm{w}}_{\alpha}\odot{\bm{g}}({\bm{\theta}}_{\alpha})\,,$ (13)
Notice that since that we have assumed $u_{\alpha,i}(0)>|v_{\alpha,i}(0)|$ for each $i\in[p]$ we have ${\bm{w}}_{\alpha}(0)>0$ entrywise. So, by (13) for all $t>0$ ,
${\bm{w}}_{\alpha}(t)>0\,.$
It suffices to track ${\bm{w}}_{\alpha}(t)$ because we can relate the log-scale magnitude of ${\bm{w}}_{\alpha}(t)$ to the magnitudes of the corresponding coordinates in ${\bm{u}}_{\alpha}(t)$ and ${\bm{v}}_{\alpha}(t)$ – see technical Lemmas D.1 D.2 and D.3.
C.5 Claimed invariants in proof of Theorem C.4
In order to prove Theorem C.4, we consider any gradient flow ${\bm{\theta}}_{\alpha}:[0,T^{*}]\to\mathbb{R}^{p}$ solving (12) where $T^{*}\in(T_{K},T_{K+1})$. For now, we focus only on proving properties of this gradient flow, and defer its existence and uniqueness to Section C.8.
We show the following invariants inductively on the stage $k$. For any $\epsilon>0$, any stage $k\leq K$, there is $\alpha_{k}:=\alpha_{k}(\epsilon)>0$ such that for all $\alpha<\alpha_{k}$ the following holds. There are times $\bar{t}_{k}:=\bar{t}_{k}(\alpha,\epsilon)$ and $\underline{t}_{k+1}:=\underline{t}_{k+1}(\alpha,\epsilon)$, such that
$\bar{t}_{k}$ $\in[T_{k}-\epsilon,T_{k}+\epsilon]\,,$ (14)
$\underline{t}_{k+1}$$\in\begin{cases}[T_{k+1}-\epsilon,T_{k+1}+\epsilon]\,,&\mbox{ if }T_{k+1}<\infty\\ \{T^{*}\},&\mbox{ if }T_{k+1}=\infty\end{cases}\,.$ (15)
and the weights approximate the greedy limit for all times $t\in[\bar{t}_{k},\underline{t}_{k+1}]$
$\|{\bm{\theta}}_{\alpha}(t)-{\bm{\theta}}^{k}\|<\epsilon\,,$ (16)
and the weights at times $\bar{t}_{k}$ and $\underline{t}_{k+1}$ are correctly estimated by the incremental learning dynamics on the logarithmic-scale
$\|\log_{\alpha}({\bm{w}}_{\alpha}(\bar{t}_{k}))-{\bm{b}}^{k}\|<\epsilon\,$ (17)
and if $T_{k+1}<\infty$ then
$\|\log_{\alpha}({\bm{w}}_{\alpha}(\underline{t}_{k+1}))-{\bm{b}}^{k+1}\|<\epsilon\,.$ (18)
Base case $k=0$: Take $\bar{t}_{0}(\alpha,\epsilon)=0$. Then statement (14) holds since $T_{0}=0$. Notice that as $\alpha\to 0$ we have that ${\bm{u}}_{\alpha}(0),{\bm{v}}_{\alpha}(0)\to{\bm{0}}={\bm{u}}^{0}$, and also $\log_{\alpha}{\bm{w}}_{\alpha}(0)\to{{\bm{1}}}={\bm{b}}^{0}$. So statement (17) follows if we take $\alpha_{0}$ small enough. In Section C.6 we show how to construct time $\underline{t}_{1}$ such that (16) and (18) hold.
Inductive step: Suppose that (14), (16), (17) and (18) hold for some iteration $k<K$. We prove them for iteration $k+1$. In Section C.7 we construct time $\bar{t}_{k}$. In Section C.6 we construct time $\underline{t}_{k+1}$.
C.6 Dynamics from time $\bar{t}_{k}$ to time $\underline{t}_{k+1}$ (Linear dynamics for $O(\log(1/\alpha))$ unrescaled time)
Let $k\leq K$, and suppose that we know that for any $\bar{\epsilon}_{k}>0$, there is $\bar{\alpha}_{k}(\bar{\epsilon}_{k})>0$ such that for all $0<\alpha<\bar{\alpha}_{k}$, there is a time $\bar{t}_{k}=\bar{t}_{k}(\alpha,\bar{\epsilon}_{k})$ satisfying
$|T_{k}-\bar{t}_{k}|$ $<\bar{\epsilon}_{k}$
$\|{\bm{\theta}}_{\alpha}(\bar{t}_{k})-{\bm{\theta}}^{k}\|$ $<\bar{\epsilon}_{k}$
$\|\log_{\alpha}({\bm{w}}_{\alpha}(\bar{t}_{k}))-{\bm{b}}^{k}\|$ $<\bar{\epsilon}_{k}\,.$
C.6.1 Analysis in case where $T_{k+1}<\infty$
Consider first the case where $T_{k+1}<\infty$. We show that, for any $\underline{\epsilon}_{k+1}>0$, there is $\rho_{k+1}(\underline{\epsilon}_{k+1})>0$ such that for all $0<\rho<\rho_{k+1}(\bar{\epsilon}_{k+1})$ there is $\underline{\alpha}_{k+1}(\rho,\underline{\epsilon}_{k+1})>0$ such that for all $\alpha<\underline{\alpha}_{k+1}$, there is a time $\underline{t}_{k+1}=\underline{t}_{k+1}(\alpha,\rho,\underline{\epsilon}_{k+1})$ satisfying
$|T_{k+1}-\underline{t}_{k+1}|$ $<\underline{\epsilon}_{k+1}$ (19)
$\|{\bm{\theta}}_{\alpha}(t)-{\bm{\theta}}^{k}\|$ $<\underline{\epsilon}_{k+1}\mbox{ for all }t\in[\bar{t}_{k},\underline{t}_{k+1}]$ (20)
$\|\log_{\alpha}({\bm{w}}_{\alpha}(\underline{t}_{k+1}))-{\bm{b}}^{k+1}\|$ $<\underline{\epsilon}_{k+1}$ (21)
$u_{\alpha,i_{k}}(\underline{t}_{k+1})$ $\in[\rho,3\rho]\,,$ (22)
$\operatorname{sgn}(v_{\alpha,i_{k}}(\underline{t}_{k+1}))$ $=s_{i_{k}}^{k+1}\,.$ (23)
For any $\rho,\alpha$, let $\bar{\epsilon}_{k}=\rho\underline{\epsilon}_{k+1}/(4p)$ and choose $\bar{t}_{k}=\bar{t}_{k}(\alpha,\bar{\epsilon}_{k})$. Then define
$\underline{t}_{k+1}$ $=\underline{t}_{k+1}(\alpha,\rho,\underline{\epsilon}_{k+1})$ (24)
$=\inf\{t\in[\bar{t}_{k},\infty):\|{\bm{u}}_{\alpha,S_{k}^{c}}(t)-{\bm{u}}_{\alpha,S_{k}^{c}}(\bar{t}_{k})\|+\|{\bm{v}}_{\alpha,S_{k}^{c}}(t)-{\bm{v}}_{\alpha,S_{k}^{c}}(\bar{t}_{k})\|>4\rho\}\,.$
Now we show that the weights ${\bm{\theta}}_{\alpha}(t)$ cannot move much from time $\bar{t}_{k}$ to $\underline{t}_{k+1}$. The argument uses the local Lipschitzness of the loss $\mathcal{L}$ (from technical Lemma D.7), and the strictness of ${\bm{\theta}}^{k}$ as a stationary point (from Assumption 4.3).
Lemma C.5 (Stability of active variables during part (A) of dynamics).
There is $\rho_{k+1}$ small enough and $\underline{\alpha}_{k+1}(\rho)$ small enough depending on $\rho$,such that for all $\rho<\rho_{k+1}$ and $\alpha<\underline{\alpha}_{k+1}$ and all $t\in[\bar{t}_{k},\underline{t}_{k+1})$,
$\|{\bm{\theta}}_{\alpha}(t)-{\bm{\theta}}^{k}\|<\rho^{\prime}:=\max(24\rho,18\sqrt{\rho K_{R_{k}}/c_{k}})\,.$ (25)
where $c_{k}$ is the strict-minimum constant from Assumption 4.3 and $K_{R_{k}}$ is the Lipschitzness constant from Lemma D.7 for the ball of radius $R_{k}=\|{\bm{\theta}}^{k}\|+1$.
Proof.
Assume by contradiction that (25) is violated at some time $t<\underline{t}_{k+1}$. Let us choose the first such time
$t^{*}=\inf\{t\in[\bar{t}_{k},\underline{t}_{k+1}):\|{\bm{u}}_{\alpha}(t^{*})-{\bm{u}}^{k}\|+\|{\bm{v}}_{\alpha}(t^{*})-{\bm{s}}^{k}\odot{\bm{u}}^{k}\|\geq\rho^{\prime}\}\,.$
Define $\tilde{{\bm{\theta}}}=(\tilde{{\bm{u}}},\tilde{{\bm{v}}})$ by
$\tilde{u}_{i}=\begin{cases}u_{\alpha,i}(t^{*}),&i\in S_{k}\\ 0,&i\not\in S_{k}\end{cases}\quad\mbox{ and }\quad\tilde{v}_{i}=\begin{cases}v_{\alpha,i}(t^{*}),&i\in S_{k}\\ 0,&i\not\in S_{k}\end{cases}\,.$
By the definition of $\underline{t}_{k+1}$, this satisfies
$\|\tilde{{\bm{u}}}-{\bm{u}}_{\alpha}(t^{*})\|$ $=\|{\bm{u}}_{\alpha,S_{k}^{c}}(t^{*})\|\leq 4\rho+\|{\bm{u}}_{\alpha,S_{k}^{c}}(\bar{t}_{k})\|\leq 4\rho+\underline{\epsilon}_{k}<5\rho\,,$
$\|\tilde{{\bm{v}}}-{\bm{v}}_{\alpha}(t^{*})\|$ $=\|{\bm{v}}_{\alpha,S_{k}^{c}}(t^{*})\|\leq 4\rho+\|{\bm{v}}_{\alpha,S_{k}^{c}}(\bar{t}_{k})\|\leq 4\rho+\underline{\epsilon}_{k}<5\rho\,.$
Also
$\|\tilde{\bm{u}}-{\bm{u}}^{k}\|+\|\tilde{\bm{v}}-{\bm{s}}^{k}\odot{\bm{u}}^{k}\|$ $=\|{\bm{u}}_{\alpha,S_{k}}(t^{*})-{\bm{z}}_{S_{k}}^{k}\|+\|{\bm{v}}_{\alpha,S_{k}}(t^{*})-{\bm{s}}^{k}_{S_{k}}\odot{\bm{z}}_{S_{k}}^{k}\|\geq\rho^{\prime}-10\rho\geq\rho^{\prime}/2\,.$
Using (a) the strict minimum Assumption 4.3 with constant $c_{k}$, since $\|\tilde{{\bm{\theta}}}-{\bm{\theta}}^{k}\|\leq\rho^{\prime}$ and we take $\rho^{\prime}$ small enough,
$\mathcal{L}({\bm{\theta}}_{\alpha}(t^{*}))$ $\geq\mathcal{L}(\tilde{{\bm{\theta}}})-4\rho K_{R_{k}}\stackrel{{\scriptstyle(a)}}{{\geq}}\mathcal{L}({\bm{\theta}}^{k})-4\rho K_{R_{k}}+\frac{c_{k}(\rho^{\prime})^{2}}{16}$
$\geq\mathcal{L}({\bm{\theta}}_{\alpha}(\bar{t}_{k}))-(4\rho+\bar{\epsilon}_{k})K_{R_{k}}+\frac{c_{k}(\rho^{\prime})^{2}}{16}>\mathcal{L}({\bm{\theta}}_{\alpha}(\bar{t}_{k}))\,.$
This is a contradiction because $\mathcal{L}$ is nondecreasing along the gradient flow. ∎
Lemma C.6 (Log-scale approximation is correct during part (A)).
There are functions $\rho_{k+1}(\underline{\epsilon}_{k+1})>0$ and $\underline{\alpha}_{k+1}(\rho,\underline{\epsilon}_{k+1})>0$ such that for all $\rho<\rho_{k+1}$ and $\alpha<\underline{\alpha}_{k+1}$, and for all $t\in(\bar{t}_{k},\underline{t}_{k+1})$ we have for a constant $C$ depending on $k$,
$\|\log_{\alpha}({\bm{w}}_{\alpha}(t))-{\bm{b}}^{k}+(t-\bar{t}_{k}){\bm{g}}({\bm{\theta}}^{k})\|<\rho\underline{\epsilon}_{k+1}+C\rho^{\prime}(t-\bar{t}_{k})\,.$ (26)
Furthermore, for all $i\in S_{k}^{c}$ and $t\in(\bar{t}_{k},\underline{t}_{k+1})$ we have
$\operatorname{sgn}(g_{i}({\bm{\theta}}_{\alpha}(t)))=\operatorname{sgn}(g_{i}({\bm{\theta}}^{k})).$ (27)
Proof.
By Lemma C.5 and Lemma D.7, there is a constant $C$ depending on ${\bm{\theta}}^{k}$ such that for all $t\in(\bar{t}_{k},\underline{t}_{k+1})$,
$\|{\bm{g}}({\bm{\theta}}_{\alpha}(t))-{\bm{g}}({\bm{\theta}}^{k})\|\leq C\rho^{\prime}\,.$
For shorthand, write $\bar{{\bm{g}}}({\bm{\theta}}^{k})={\bm{g}}({\bm{\theta}}^{k})+C\rho^{\prime}{{\bm{1}}}$ and $\underline{{\bm{g}}}({\bm{\theta}}^{k})={\bm{g}}({\bm{\theta}}^{k})-C\rho^{\prime}{{\bm{1}}}$. Since ${\bm{w}}_{\alpha}(t)>0$ entrywise as we have assumed without loss of generality (see Section C.3), we have the following entrywise inequalities
$\underline{{\bm{g}}}({\bm{\theta}}^{k})\odot{\bm{w}}_{\alpha}(t)<{\bm{g}}({\bm{\theta}}_{\alpha}(t))\odot{\bm{w}}_{\alpha}(t)<\bar{{\bm{g}}}({\bm{\theta}}^{k})\odot{\bm{w}}_{\alpha}(t)\,.$ (28)
Since the dynamics are given by $\frac{d{\bm{w}}_{\alpha}}{dt}=\log(1/\alpha){\bm{g}}({\bm{w}}_{\alpha})\odot{\bm{w}}_{\alpha}$,
${\bm{w}}_{\alpha}(\bar{t}_{k})e^{(t-\bar{t}_{k})\log(1/\alpha)\underline{{\bm{g}}}({\bm{\theta}}^{k})}\leq{\bm{w}}_{\alpha}(t)\leq{\bm{w}}_{\alpha}(\bar{t}_{k})e^{(t-\bar{t}_{k})\log(1/\alpha)\bar{{\bm{g}}}({\bm{\theta}}^{k})}\,.$
Taking the logarithms with base $\alpha\in(0,1)$,
$(t-\bar{t}_{k})\underline{{\bm{g}}}({\bm{u}}^{k})\leq\log_{\alpha}({\bm{w}}_{\alpha}(\bar{t}_{k}))-\log_{\alpha}({\bm{w}}_{\alpha}(t))\leq(t-\bar{t}_{k})\bar{{\bm{g}}}({\bm{u}}^{k})\,.$
The bound (26) follows since $\|\log_{\alpha}({\bm{w}}_{\alpha}(\bar{t}_{k}))-{\bm{b}}^{k}\|<\bar{\epsilon}_{k}<\rho\underline{\epsilon}_{k+1}$.
Finally, the claim (27) follows from (28) since $\operatorname{sgn}(\bar{{\bm{g}}}({\bm{\theta}}^{k}))=\operatorname{sgn}(\underline{{\bm{g}}}({\bm{\theta}}^{k}))=\operatorname{sgn}({\bm{g}}({\bm{\theta}}^{k}))$ if we take $\rho$ small enough. ∎
First, we show that the weights must move significantly by time roughly $T_{k+1}$. This is because of the contribution of coordinate $i_{k}$.
Lemma C.7 ($\underline{t}_{k+1}$ is not much larger than $T_{k+1}$).
Suppose that $T_{k+1}<\infty$. Then there are $\rho_{k+1}(\underline{\epsilon}_{k+1})>0$ and $\underline{\alpha}_{k+1}(\rho,\underline{\epsilon}_{k+1})>0$ such that for all $\rho<\rho_{k+1}$ and $\alpha<\underline{\alpha}_{k+1}$, the following holds.
$\underline{t}_{k+1}<T_{k+1}+\underline{\epsilon}_{k+1}\,.$
Proof.
Assume by contradiction that $\underline{t}_{k+1}<T_{k+1}+\underline{\epsilon}_{k+1}$. For all times $t\in[\bar{t}_{k},\min(\underline{t}_{k+1},T_{k+1}+\underline{\epsilon}_{k+1})]$, by Lemma C.6,
$|\log_{\alpha}(w_{\alpha,i_{k}}(t))-b_{i_{k}}^{t}+(t-\bar{t}_{k})g_{i_{k}}({\bm{\theta}}^{k})|<O(\sqrt{\rho})\,.$
Since we know $|\Delta_{k}(i_{k})-(T_{k+1}-\bar{t}_{k})|<\bar{\epsilon}_{k}$ and $b_{i}^{k}-\Delta_{k}(i_{k})g_{i_{k}}({\bm{\theta}}^{k})\in\{0,2\}$, it follows that
$\log_{\alpha}(w_{\alpha,i_{k}}(T_{k+1}+\underline{\epsilon}_{k+1}))\not\in(-|g_{i_{k}}({\bm{\theta}}^{k})|(\underline{\epsilon}_{k+1}-\bar{\epsilon}_{k+1}),2+|g_{i_{k}}({\bm{\theta}}^{k})|(\underline{\epsilon}_{k+1}-\bar{\epsilon}_{k+1}))+O(\sqrt{\rho}).$
By taking $\rho$ small enough, we see that $|g_{i_{k}}({\bm{\theta}}^{k})|(\underline{\epsilon}_{k+1}-\bar{\epsilon}_{k+1})+O(\sqrt{\rho})>\delta>0$ for some $\delta>0$ that is independent of $\alpha$, so
$\log_{\alpha}(w_{\alpha,i_{k}}(T_{k+1}+\underline{\epsilon}_{k+1}))\not\in(-\delta,2+\delta)\,.$
So $|u_{\alpha,i_{k}}(T_{k+1}+\underline{\epsilon}_{k+1})|>1$ by Lemma D.2. But by the construction of $\underline{t}_{k+1}$ this means that $\underline{t}_{k+1}<T_{k+1}+\underline{\epsilon}_{k+1}$. ∎
Next, we show that until time $\underline{t}_{k+1}$, none of the coordinates in $S_{k}^{c}$ move significantly, with the possible exception of coordinate $i_{k}$.
Lemma C.8 (No coordinates in $S_{k}^{c}\setminus\{i_{k}\}$ move significantly during part (A)).
Suppose $T_{k+1}<\infty$. Then there are $\rho_{k+1}(\underline{\epsilon}_{k+1})>0$ and $\underline{\alpha}_{k+1}(\rho,\underline{\epsilon}_{k+1})>0$ such that for all $\rho<\rho_{k+1}$ and $\alpha<\underline{\alpha}_{k+1}$, the following holds. There is a constant $c>0$ depending on $k$ such that for all $i\in S_{k}^{c}\setminus\{i_{k}\}$ and $t\in[\bar{t}_{k},\underline{t}_{k+1}]$,
$|u_{\alpha,i}(t)-u_{\alpha,i}(\bar{t}_{k})|,|v_{\alpha,i}(t)-v_{\alpha,i}(\bar{t}_{k})|<\alpha^{c}+\bar{\epsilon}_{k}\,.$
Proof.
The previous lemma combined with the inductive hypothesis gives
$\underline{t}_{k+1}-\bar{t}_{k}<\Delta_{k}(i_{k})+2\underline{\epsilon}_{k+1}\setminus\{i_{k}\}.$
We analyze the movement of each coordinate $i\in S_{k}^{c}\setminus\{i_{k}\}$ by breaking into two cases:
  • Coordinate $i\neq i_{k}$ such that $b_{i}^{k}\in(0,2)$. By Assumption 4.2, there is a unique winning coordinate so $b_{i}^{k}-\tau g_{i}({\bm{\theta}}^{k})\in(c,2-c)$ for some constant $c>0$ for all $\tau\in[0,\underline{t}_{k+1}-\bar{t}_{k}]\subseteq[0,\Delta_{k}(i_{k})+2\underline{\epsilon}_{k+1}]$. By Lemma C.6, $\log_{\alpha}(w_{\alpha,i}(t))\in(-c/2,2-c/2)$ for all times $t\in[\bar{t}_{k},\underline{t}_{k+1}]$. So by Lemma D.1, $|u_{\alpha,i}(t)|,|v_{\alpha,i}(t)|\leq\alpha^{c/4}$.
  • Coordinate $i\neq i_{k}$ such that $b_{i}^{k}=0$. By Lemma D.4, we must be in the corner case where $i\in S_{k-1}\cap S_{k}^{c}$ (i.e., the coordinate was active in the previous stage but was dropped from the support in this stage).
    By Lemma D.4, since $b_{i}^{k}=0$ we have $g_{i}({\bm{\theta}}^{k})<0$. By Lemma C.6, this means $\operatorname{sgn}(g_{i}({\bm{\theta}}_{\alpha}(t)))=\operatorname{sgn}(g_{i}({\bm{\theta}}^{k}))<0$ for all $t\in(\bar{t}_{k},\underline{t}_{k+1})$.
    We break the analysis into two parts. Since $b_{i}^{k}=0$, the sign is $s_{i}^{k}=+1$. The inductive hypothesis $\|{\bm{\theta}}_{\alpha}(\bar{t}_{k})-{\bm{\theta}}^{k}\|<\bar{\epsilon}_{k}$ implies that $|u_{\alpha,i}(\bar{t}_{k})-z_{i}^{k}|<\bar{\epsilon}_{k}$ and $|v_{\alpha,i}(\bar{t}_{k})-z_{i}^{k}|<\bar{\epsilon}_{k}$. For small enough $\bar{\epsilon}_{k}$ this means that $\operatorname{sgn}(u_{\alpha,i}(\bar{t}_{k}))=\operatorname{sgn}(v_{\alpha,i}(\bar{t}_{k}))=+1$. Now let $t^{*}=\min(\underline{t}_{k+1},\inf\{t>\bar{t}_{k}:v_{\alpha,i}(t)=0\})$. Since $u_{\alpha,i}(t)>v_{\alpha,i}(t)$ without loss of generality (see Section C.3), we have $\operatorname{sgn}(u_{\alpha,i}(t))=\operatorname{sgn}(v_{\alpha,i}(t))=+1$ for all $t\in[\bar{t}_{k},t^{*}]$. So $\frac{du_{\alpha,i}(t)}{dt},\frac{dv_{\alpha,i}(t)}{dt}<0$ for all $t\in[\bar{t}_{k},t^{*}]$. So, for any $t\in[\bar{t}_{k},t^{*}]$,
    $|u_{\alpha,i}(t)-u_{\alpha,i}(\bar{t}_{k})|,|v_{\alpha,i}(t)-v_{\alpha,i}(\bar{t}_{k})|<\bar{\epsilon}_{k}$
    Also, since $\log_{\alpha}(w_{\alpha,i}(t^{*}))\approx 1$, by Lemma C.6 we have $t^{*}>c>0$ for some constant $c$ independent of $\alpha$. So for all $t\in[t^{*},\underline{t}_{k+1}]$ we have $b_{i}^{k}-\tau g_{i}({\bm{\theta}}^{k})\in(c,2-c)$ for some constant $c>0$. So $|u_{\alpha,i}(t)|,|v_{\alpha,i}(t)|\leq\alpha^{c/4}$ for all $t\in[t^{*},\underline{t}_{k+1}]$. The conclusion follows by triangle inequality.
  • Coordinate $i\neq i_{k}$ such that $b_{i}^{k}=2$. The analysis is analogous to the case $b_{i}^{k}=0$, except that we have $s_{i}^{k}=-1$ instead and $g_{i}({\bm{\theta}}^{k})>0$ by Lemma D.4.
Finally, we use this conclude that $\underline{t}_{k+1}\approx T_{k+1}$ and that the weights at coordinate $i_{k}$ are the only weights that change significantly, and by an amount approximately $\rho$.
Lemma C.9 (Coordinate $i_{k}$ wins the part (A) race at time $\underline{t}_{k+1}\approx T_{k+1}$).
Suppose that $T_{k+1}<\infty$. Then there are $\rho_{k+1}(\underline{\epsilon}_{k+1})>0$ and $\underline{\alpha}_{k+1}(\rho,\underline{\epsilon}_{k+1})>0$ such that for all $\rho<\rho_{k+1}$ and $\alpha<\underline{\alpha}_{k+1}$, the following holds.
$|\underline{t}_{k+1}-T_{k+1}|<\underline{\epsilon}_{k+1}\,,$
$u_{\alpha,i_{k}}(\underline{t}_{k+1})\in[\rho,3\rho]\,,$
$\operatorname{sgn}(v_{\alpha,i_{k}}(\underline{t}_{k+1}))=s_{i_{k}}^{k+1}\,.$
Proof.
Let us analyze the case that $b_{i_{k}}^{k}\in(0,2)$. Notice that $b_{i_{k}}^{k+1}=b_{i_{k}}^{k}-\Delta_{k}(i_{k})g_{i_{k}}({\bm{\theta}}^{k})\in\{0,2\}$ and that if $b_{i}^{k+1}=0$ then $g_{i_{k}}({\bm{\theta}}^{k})>0$ and if it is $2$ then $b_{i_{k}}^{k+1}=g_{i_{k}}({\bm{\theta}}^{k})<0$. So by Lemma C.6, for all times $t\in[\bar{t}_{k},\min(\underline{t}_{k+1},T_{k+1}-\underline{\epsilon}_{k+1})]$, we have $w_{\alpha,i_{k}}(t)\in(c,2-c)$ for some $c>0$. So for small enough $\alpha$ by Lemma D.1, $|u_{\alpha,i_{k}}(t)|,|v_{\alpha,i_{k}}(t)|\leq\alpha^{c/2}$. Combining this with Lemma C.8, we see that for $t\in[\bar{t}_{k},\min(\underline{t}_{k+1},T_{k+1}-\underline{\epsilon}_{k+1})]$ we have
$\|{\bm{u}}_{\alpha}(t)-{\bm{u}}_{\alpha}(\bar{t}_{k})\|+\|{\bm{v}}_{\alpha}(t)-{\bm{v}}_{\alpha}(\bar{t}_{k})\|<2(\alpha^{c}+\bar{\epsilon}_{k})p<\rho\,,$
for small enough $\alpha$. So by definition of $\underline{t}_{k+1}$ we must have $\underline{t}_{k+1}>T_{k+1}-\underline{\epsilon}_{k+1}$. Combined with Lemma C.7, we conclude that $|T_{k+1}-\underline{t}_{k+1}|<\underline{\epsilon}_{k+1}$, which is the first claim of the lemma. Furthermore, by Lemma C.8,
$\sum_{i\in S_{k}^{c}\setminus\{i_{k}\}}|u_{\alpha,i}(\underline{t}_{k+1})-u_{\alpha,i}(\bar{t}_{k})|+|v_{\alpha,i}(\underline{t}_{k+1})-v_{\alpha,i}(\bar{t}_{k})|\leq 2p(\alpha^{c}+\bar{\epsilon}_{k}))<\rho/2,$
so by definition of $\underline{t}_{k+1}$ and triangle inequality we have $|u_{\alpha,i_{k}}(\underline{t}_{k+1})|+|v_{\alpha,i_{k}}(\underline{t}_{k+1})|\geq 4\rho-\rho/2=7\rho/2$. Also, since $u_{\alpha,i_{k}}^{2}(\underline{t}_{k+1})-v_{\alpha,i_{k}}^{2}(\underline{t}_{k+1})=\Theta(\alpha^{2})$ we have $u_{\alpha,i_{k}}(\underline{t}_{k+1})\in[\rho,3\rho]$. Finally, if $b_{i_{k}}^{k+1}=2$, then $s_{i_{k}}^{k+1}=-1$ and $\log_{\alpha}(w_{\alpha,i_{k}}(\underline{t}_{k+1}))>1.5$ so $\operatorname{sgn}(v_{\alpha,i_{k}}(t))<0$ by Lemma D.3; analogously, if $b_{i_{k}}^{k+1}=0$, we have $s_{i_{k}}^{k+1}=1$ and $\log_{\alpha}(w_{\alpha,i_{k}}(\underline{t}_{k+1})<0.5$ so $\operatorname{sgn}(v_{\alpha,i_{k}}(\underline{t}_{k+1})>0$.
The case $b_{i_{k}}^{k}\in\{0,2\}$ can be proved similarly to the analysis in Lemma C.8, where one shows that during the first period of time the magnitudes of $|u_{i_{k}}(t)|$ and $|v_{i_{k}}(t)|$ decrease, until the sign of $v_{i_{k}}$ flips and they once again increase.
We have shown the claims (19), (20), (21) (22), and (23) for the time $\underline{t}_{k+1}$. In fact, if we let $\underline{t}_{k+1}^{\prime}\in[\bar{t}_{k},\infty)$ be the first time $t$ such that $u_{\alpha,i_{k}}(t)=\rho$ we still have (19), (20), (21) and (23) by the same analysis as above, and (22) can be replaced with the slightly more convenient
$u_{\alpha,i_{k}}(\underline{t}_{k+1}^{\prime})=\rho\,.$
C.6.2 Analysis in case where $T_{k+1}=\infty$
In this case that $T_{k+1}$, we just have to show that the weights remain close to ${\bm{\theta}}^{k}$. We show that for any $\underline{\epsilon}_{k+1}>0$, there is $\underline{\alpha}_{k+1}(\underline{\epsilon}_{k+1})>0$ such that for all $\alpha<\underline{\alpha}_{k+1}$ and times $t\in[T_{k}+\underline{\epsilon}_{k+1},T^{*}]$,
$\|{\bm{\theta}}_{\alpha}(t)-{\bm{\theta}}^{k}\|<\underline{\epsilon}_{k+1}.$
We can use Lemmas C.5 and C.6, which were developed for the case of $T_{k+1}<\infty$, but still hold for $T_{k+1}=\infty$. Lemma C.5 guarantees that the weights do not move much until time $\underline{t}_{k+1}$, and so we only need to show that $\underline{t}_{k+1}\geq T^{*}$ when we take $\rho$ small enough. For this, observe that $g_{i}({\bm{\theta}}^{k})=0$ for all $i\not\in S_{k}$, because otherwise $T_{k+1}<\infty$. Therefore Lemma C.6 guarantees that until time $\min(T_{*},\underline{t}_{k+1})$ all weights are close to the original on the logarithmic scale. Namely,
$\|\log_{\alpha}({\bm{w}}_{\alpha}(t))-{\bm{b}}^{k}\|<\rho\underline{\epsilon}_{k+1}+C\rho^{\prime}(T^{*}-\bar{t}_{k})$
Furthermore, by the non-degeneracy Assumption 4.2 we know that $b_{i}^{k}\in(0,2)$ for all $i\not\in S_{k}$ by Lemma D.4. So if we take $\rho$ small enough and $\underline{\alpha}_{k+1}$ small enough, we must have that $\underline{t}_{k+1}\geq T^{*}$.
C.7 Dynamics from time $\underline{t}_{k}$ to time $\bar{t}_{k}$ (Nonlinear evolution for $O(1)$ unrescaled time)
Suppose that we know for some $k\leq K$ that for any $\underline{\epsilon}_{k}>0$, there is $\rho_{k}(\underline{\epsilon}_{k})>0$ such that for all $\rho<\rho_{k}$ there is $\underline{\alpha}_{k}(\rho,\underline{\epsilon}_{k})>0$ such that for all $\alpha<\underline{\alpha}_{k}$, there is a time $\underline{t}_{k}=\underline{t}_{k}(\alpha,\rho,\underline{\epsilon}_{k})$ satisfying
$|T_{k}-\underline{t}_{k}|$ $<\underline{\epsilon}_{k}$ (29)
$\|{\bm{\theta}}_{\alpha}(\underline{t}_{k})-{\bm{\theta}}^{k-1}\|$ $<\underline{\epsilon}_{k}$ (30)
$\|\log_{\alpha}({\bm{w}}_{\alpha}(\underline{t}_{k}))-{\bm{b}}^{k}\|$ $<\underline{\epsilon}_{k}$ (31)
$u_{\alpha,i_{k-1}}(\underline{t}_{k})$ $=\rho\,,$ (32)
$\operatorname{sgn}(v_{\alpha,i_{k-1}}(\underline{t}_{k}))$ $=s_{i_{k-1}}^{k}\,.$ (33)
Now we will show that for any $\bar{\epsilon}_{k}>0$, there is $\bar{\alpha}_{k}=\bar{\alpha}_{k}(\bar{\epsilon}_{k})>0$ such that for all $0<\alpha<\bar{\alpha}_{k}$, there is a time $\bar{t}_{k}=\bar{t}_{k}(\alpha,\bar{\epsilon}_{k})$ satisfying
$|T_{k}-\bar{t}_{k}|<\bar{\epsilon}_{k}$ (34)
$\|{\bm{\theta}}_{\alpha}(\bar{t}_{k})-{\bm{\theta}}^{k}\|<\bar{\epsilon}_{k}$ (35)
$\|\log_{\alpha}({\bm{w}}_{\alpha}(\bar{t}_{k}))-{\bm{b}}^{k}\|<\bar{\epsilon}_{k}$ (36)
We give the construction for $\bar{t}_{k}$. For any desired accuracy $\bar{\epsilon}_{k}>0$ in this stage, we will construct an accuracy $\underline{\epsilon}_{k}=\underline{\epsilon}_{k}(\bar{\epsilon}_{k})=\bar{\epsilon}_{k}/3>0$. We will also construct a $\rho=\rho(\underline{\epsilon}_{k})>0$ which is sufficiently small, and we will construct an cutoff for $\alpha$ equal to $\bar{\alpha}_{k}=\bar{\alpha}_{{k+1}}(\bar{\epsilon}_{k})>0$ which satisfies $\bar{\alpha}_{k}<\underline{\alpha}_{k}(\rho,\underline{\epsilon}_{k})$. The values for these parameters $\underline{\epsilon}_{k}$ and $\rho$ and $\bar{\alpha}_{k}$ will be chosen in the following lemma, and will depend only on $\bar{\epsilon}_{k}$.
Lemma C.10 (New local minimum reached in time $O(1/\log(1/\alpha))$).
For any $\bar{\epsilon}_{k}>0$, we can choose $\bar{\alpha}_{k}=\bar{\alpha}_{k}(\bar{\epsilon}_{k})>0$ small enough so that, for any $0<\alpha<\bar{\alpha}_{k}$, there is $\bar{t}_{k}=\bar{t}_{k}(\alpha,\bar{\epsilon}_{k})$ for which conditions (34) to (36) hold.
Furthermore, there is a constant $C^{\prime\prime}$ independent of $\alpha$ such that $|{\bm{\theta}}_{\alpha}(t)|/|{\bm{\theta}}_{\alpha}(\underline{t}_{k})|\in[1/C^{\prime\prime},C^{\prime\prime}]^{2p}$ at all times $t\in[\underline{t}_{k},\bar{t}_{k}]$.
Proof.
Let $\underline{t}_{k}=\underline{t}_{k}(\alpha,\rho,\underline{\epsilon}_{k})$ be given by the induction. Let us compare the dynamics starting at ${\bm{\theta}}_{\alpha}(\underline{t}_{k})$ with the dynamics starting at $\tilde{{\bm{\theta}}}(\underline{t}_{k})=(\tilde{{\bm{u}}}(\underline{t}_{k}),\tilde{{\bm{v}}}(\underline{t}_{k}))$ which is given by
$\tilde{u}_{i}(\underline{t}_{k})=\begin{cases}u_{\alpha,i}(\underline{t}_{k}),&i\in S_{k-1}\cup\{i_{k-1}\}\\ 0,&\mbox{otherwise}\end{cases}\quad\mbox{ and }\quad\tilde{v}_{i}(\underline{t}_{k})=\begin{cases}v_{\alpha,i}(\underline{t}_{k}),&i\in S_{k-1}\cup\{i_{k-1}\}\\ 0,&\mbox{otherwise}\end{cases}$
and run with
$\frac{d\tilde{{\bm{\theta}}}}{dt}=-\log(1/\alpha)\nabla_{{\bm{w}}}\mathcal{L}(\tilde{{\bm{\theta}}})\,\,.$
By Assumption 4.4 we know there exists a unique solution $\tilde{{\bm{\theta}}}:[\underline{t}_{k},\infty)\to\mathbb{R}^{p}$ as long as we take $\underline{\epsilon}_{k}$ small enough because $\mathrm{supp}(\tilde{{\bm{\theta}}}(\underline{t}_{k}))=S_{k-1}\cup\{i_{k-1}\}$ and $\|\tilde{{\bm{\theta}}}_{i}(\underline{t}_{k})-{\bm{\theta}}^{k-1}\|<\underline{\epsilon}_{k}$. Furthermore, by Assumption 4.4 if we take $\underline{\epsilon}_{k}$ small enough there must be a time $\tau:=\tau(\bar{\epsilon}_{k},\rho)<\infty$ such that
$\|\tilde{{\bm{\theta}}}(t)-{\bm{\theta}}^{k}\|<\bar{\epsilon}_{k}/2\mbox{ for }t\geq\underline{t}_{k}+\tau/\log(1/\alpha)$ (37)
Define
$\bar{t}_{k}=\underline{t}_{k}+\tau/\log(1/\alpha).$
So for $\alpha$ small enough, $|T_{k}-\bar{t}_{k}|<2\underline{\epsilon}_{k}<\bar{\epsilon}_{k}$, proving (34).
We now compare ${\bm{\theta}}_{\alpha}(\bar{t}_{k})$ with $\tilde{{\bm{\theta}}}(\bar{t}_{k})$, and show that if we take $\alpha$ small enough, then the dynamics of $\tilde{{\bm{\theta}}}$ closely match the dynamics of ${\bm{\theta}}_{\alpha}(t)$ for times $\underline{t}_{k}+O(1/\log(1/\alpha))$. The argument uses Gronwall’s inequality. Let $t^{*}=\inf\{t>\underline{t}_{k}:\|\tilde{{\bm{\theta}}}(t^{*})-{\bm{\theta}}_{\alpha}(t)\|>1/3\}$. For times $t\in[\underline{t}_{k},t^{*})$ by Lemma D.7 we have
$\|\frac{d}{dt}\tilde{{\bm{\theta}}}(t)-\frac{d}{dt}{\bm{\theta}}_{\alpha}(t)\|$ $=\log(1/\alpha)\|\nabla_{{\bm{\theta}}}\mathcal{L}(\tilde{{\bm{\theta}}}(t))-\nabla_{{\bm{\theta}}}\mathcal{L}({\bm{\theta}}_{\alpha}(t))\|\leq K_{\tilde{{\bm{\theta}}}(t)}\log(1/\alpha)\|\tilde{{\bm{\theta}}}(t)-{\bm{\theta}}_{\alpha}(t)\|,$
where $K_{\tilde{{\bm{\theta}}}(t)}$ is the smoothness constant from Lemma D.7. Note that since $\|\tilde{{\bm{\theta}}}(t)\|<\infty$ for large enough $t$ by (37), the trajectory of $\tilde{{\bm{\theta}}}$ must lie in a compact set. Therefore, there must be a finite set of times $s_{1},\ldots,s_{m}\in[\underline{t}_{k},t^{*})$ such that $\cup_{t\in[\underline{t}_{k},t^{*})}B(\tilde{{\bm{\theta}}}(t),1/2)\subseteq\cup_{i=1}^{m}B(\tilde{{\bm{\theta}}}(s_{i}),3/4)$. So letting $C=\max_{i=1}^{m}K_{\tilde{{\bm{\theta}}}(s_{i})}<\infty$ for all times $t\in[\underline{t}_{k},t^{*})$ we have
$\frac{d}{dt}\|\tilde{{\bm{\theta}}}(t)-{\bm{\theta}}_{\alpha}(t)\|\leq C\log(1/\alpha)\|\tilde{{\bm{\theta}}}(t)-{\bm{\theta}}_{\alpha}(t)\|\,.$
By Gronwall’s inequality, for all times $t\in[\underline{t}_{k},t^{*})$,
$\|\tilde{{\bm{\theta}}}(t)-{\bm{\theta}}_{\alpha}(t)\|\leq\|\tilde{{\bm{\theta}}}(\underline{t}_{k})-{\bm{\theta}}_{\alpha}(\underline{t}_{k})\|\exp(C\log(1/\alpha)(t-\underline{t}_{k}))\,.$
We know from Lemma C.8 that there is a constant $c>0$ such that for any small enough $0<\alpha<\underline{\alpha}_{k}$, such that
$\|\tilde{{\bm{\theta}}}(\underline{t}_{k})-{\bm{\theta}}_{\alpha}(\underline{t}_{k})\|<\alpha^{c}$
If we take $\alpha$ small enough that $\alpha^{c}\exp(C\tau)<\bar{\epsilon}_{k}/2<1/3$, we must have $t^{*}>\underline{t}_{k}+\tau/\log(1/\alpha)$ and so we prove (35)
$\|{\bm{\theta}}^{k}-{\bm{\theta}}_{\alpha}(\bar{t}_{k})\|\leq\bar{\epsilon}_{k}/2+\|\tilde{{\bm{\theta}}}(\bar{t}_{k})-{\bm{\theta}}_{\alpha}(\bar{t}_{k})\|<\bar{\epsilon}_{k}\,.$
It remains to show that (36) is satisfied. Since $\|\tilde{{\bm{\theta}}}(t)-{\bm{\theta}}_{\alpha}(t)\|<1/3$ for all $t\in[\underline{t}_{k},\bar{t}_{k}]$, it holds that the trajectory of ${\bm{\theta}}_{\alpha}(t)$ lies in a compact set. So by Lemma D.7 we have $\|{\bm{g}}({\bm{\theta}}_{\alpha}(t))\|<C^{\prime}$ for some constant $C^{\prime}$ at all times $t\in[\underline{t}_{k},\bar{t}_{k}]$. Since $\frac{1}{\log(1/\alpha)}|\frac{dw_{\alpha,i}}{dt}|=|w_{\alpha,i}(t)||g_{i}({\bm{w}}_{\alpha}(t))|<C^{\prime}|w_{\alpha,i}(t)|$, we must have $|w_{\alpha,i}(t)|/|w_{\alpha,i}(\underline{t}_{k})|\in[1/C^{\prime\prime},C^{\prime\prime}]$ for some constant $C^{\prime\prime}$ independent of $\alpha$ and all $t\in[\underline{t}_{k},\bar{t}_{k}]$. Therefore, (36) follows from (31). A similar argument shows that $|{\bm{\theta}}_{\alpha}(t)/{\bm{\theta}}_{\alpha}(\underline{t}_{k})|\in[1/C^{\prime\prime},C^{\prime\prime}]^{2p}$.
C.8 Concluding the proof of Theorem C.4
We have shown that Theorem 4.5 is true for solutions ${\bm{\theta}}_{\alpha}:[0,T^{*}]\to\mathbb{R}^{2p}$ to the gradient flow, where $T_{*}\in(T_{K},T_{K+1})$. To establish Theorem C.4 it remains only to show that for any $T_{*}\in(T_{K},T_{K+1})$ and small enough $\alpha$ such a solution to the gradient flow exists and is unique. To see this, note that in the inductive proof of the invariants we construct a sequence of times $0=\bar{t}_{0}\leq\underline{t}_{1}\leq\bar{t}_{1}\leq\dots\leq\bar{t}_{K}\leq\underline{t}_{K+1}>T_{*}$, where we guarantee that any gradient flow solution ${\bm{\theta}}_{\alpha}:[0,\underline{t}_{k+1}]\to\mathbb{R}^{p}$ satisfies ${\bm{\theta}}_{\alpha}\in\cup_{k\in\{0,\ldots,K\}}B({\bm{\theta}}^{k},1)$ for all $t\in\cup_{k\in\{0,\ldots,K\}}[\bar{t}_{k},\underline{t}_{k+1}]$. And also for $t\in\cup_{k\in\{0,\ldots,K-1\}}[\underline{t}_{k},\bar{t}_{k+1}]$, we have ${\bm{\theta}}_{\alpha}(t)\in B(0,C^{\prime\prime}_{k}{\bm{\theta}}^{k})$ for some constant $C^{\prime\prime}_{k}$ independent of $\alpha$ by Lemma C.10. So ${\bm{\theta}}_{\alpha}(t)\in B(0,C_{K})$ for some constant $C_{K}$ at all times $t\in[0,T^{*}]$. By Lemma D.7, the loss gradient $\nabla_{{\bm{\theta}}}\mathcal{L}({\bm{\theta}})=({\bm{v}}\odot{\bm{g}}({\bm{\theta}}),{\bm{u}}\odot{\bm{g}}({\bm{\theta}}))$ is Lipschitz-continuous on the compact set $B(0,C_{K})$. So ${\bm{\theta}}_{\alpha}:[0,T^{*}]\to\mathbb{R}^{p}$ exists and is unique by the Cauchy-Lipschitz theorem.
Appendix D Technical lemmas
D.1 Relating the sum of the weights to the original weights using the conservation law
Lemma D.1.
If for some constant $10<c<$ 1 we have $\log_{\alpha}(w_{\alpha,i}(t))\in(c,2-c)$, then for small enough $\alpha$
$\max(|u_{\alpha,i}(t)|,|v_{\alpha,i}(t)|)\leq\alpha^{c/2}\,.$
Proof.
Let $\tilde{{\bm{w}}}_{\alpha}(t)={\bm{u}}_{\alpha}(t)-{\bm{v}}_{\alpha}(t)$. By the conservation law (5), $w_{\alpha,i}(t)\tilde{w}_{\alpha,i}(t)=w_{\alpha,i}(0)\tilde{w}_{\alpha,i}(0)=u_{\alpha,i}(0)^{2}-v_{\alpha,i}(0)^{2}$. By the non-degeneracy of initialization (Assumption 4.2), the right-hand-side is $\Theta(\alpha^{2})$. So if $\log_{\alpha}(w_{\alpha,i}(t))\in(c,2-c)$ then for small enough $\alpha$, we have $\log_{\alpha}(|\tilde{w}_{\alpha,i}(t)|)\in(3c/4,2-3c/4)$. So $|u_{\alpha,i}(t)|\leq|w_{\alpha,i}(t)+\tilde{w}_{\alpha,i}(t)|\leq\alpha^{c/2}$ and $|v_{\alpha,i}(t)|\leq|w_{\alpha,i}(t)-\tilde{w}_{\alpha,i}(t)|\leq\alpha^{c/2}$. ∎
Lemma D.2.
If for some constant $0<c$ we have $\log_{\alpha}(w_{\alpha,i}(t))\not\in(-c,2+c)$, then for small enough $\alpha$,
$|u_{\alpha,i}(t)|>1\,.$
Proof.
Define $\tilde{{\bm{w}}}_{\alpha}={\bm{u}}_{\alpha}-{\bm{v}}_{\alpha}$ as in the proof of Lemma D.1. If $\log_{\alpha}(w_{\alpha,i}(t))<-c$ then $\log_{\alpha}(|\tilde{w}_{\alpha,i}(t)|)>2-c/2$ for small enough $\alpha$, so $u_{i}(t)>\alpha^{-c}-\alpha^{2-c/2}>1$. Similarly, if $\log_{\alpha}(w_{\alpha,i}(t))>2+c$ then $\log_{\alpha}(|\tilde{w}_{\alpha,i}(t)|)<-c/2$ so $|u_{i}(\alpha)|>\alpha^{-c/2}-\alpha^{2+c}>1$. ∎
Lemma D.3.
If for some constant $c>0$, there is small enough $\alpha$ such that if we have $\log_{\alpha}(w_{\alpha,i}(t))>1+c$ then $\operatorname{sgn}(v_{\alpha,i}(t))<0$. Otherwise, if $\log_{\alpha}(w_{\alpha,i}(t))<1-c$ then $\operatorname{sgn}(v_{\alpha,i}(t))>0$.
Proof.
Follows from ${\bm{v}}_{\alpha}=\frac{1}{2}({\bm{w}}_{\alpha}-\tilde{{\bm{w}}}_{\alpha})$. Recall that ${\bm{w}}_{\alpha}(t)>0$ and notice that $\tilde{{\bm{w}}}_{\alpha}(t)>0$. In the first case, $w_{\alpha,i}(t)<\alpha^{1+c}$ and $\tilde{w}_{\alpha,i}(t)>\alpha^{1-c/2}$. In the latter case $w_{\alpha,i}(t)>\alpha^{1-c}$ and $\tilde{w}_{\alpha,i}(t)<\alpha^{1+c/2}$. ∎
D.2 Sign of gradients on coordinates that leave support
Lemma D.4.
For any $k\geq 1$ and $i\in S_{k}^{c}$, if $b_{i}^{k}\in\{0,2\}$ then we must have $i\in\mathrm{supp}({\bm{u}}^{k-1})\setminus\mathrm{supp}({\bm{u}}^{k})$, and we must have $g_{i}({\bm{u}}^{k})<0$ if $b_{i}^{k}=0$ and $g_{i}({\bm{\theta}}^{k})>0$ if $b_{i}^{k}=2$. In particular, $\Delta_{k}(i_{k})>0$ for all $k$.
Proof.
This is by induction on $k$ and using the non-degeneracy Assumption 4.2. ∎
D.3 Local lipschitzness and smoothness
We provide several technical lemmas on the local Lipschitzness and smoothness of $\ell$, $h$, and ${\bm{g}}$.
Lemma D.5.
The function $\ell({\bm{y}},\cdot)$ is locally Lipschitz and smooth in its second argument: for any $R>0$, there exists $K_{R}$ such that for any ${\bm{\zeta}},{\bm{\zeta}}^{\prime}\in B(0,R)$
$|\ell({\bm{y}},{\bm{\zeta}})-\ell({\bm{y}},{\bm{\zeta}}^{\prime})|\leq K_{R}\|{\bm{\zeta}}-{\bm{\zeta}}^{\prime}\|$
$\|D\ell({\bm{y}},{\bm{\zeta}})-D\ell({\bm{y}},{\bm{\zeta}}^{\prime})\|\leq K_{R}\|{\bm{\zeta}}-{\bm{\zeta}}^{\prime}\|,$
almost surely over ${\bm{y}}$. Here $D\ell({\bm{y}},\cdot)^{\top}\in\mathbb{R}^{d_{out}}$ is the derivative in the second argument.
Proof.
Since $\ell$ is continuously twice-differentiable, for each ${\bm{y}}\in\mathbb{R}^{d_{y}},{\bm{\zeta}}\in\mathbb{R}^{d_{out}}$ there is $K_{{\bm{y}},{\bm{\zeta}}}<\infty$ such that for all ${\bm{y}}\in B({\bm{y}},1/K_{{\bm{y}},{\bm{\zeta}}})$ and ${\bm{\zeta}}^{\prime}\in B({\bm{\zeta}},1/K_{{\bm{y}},{\bm{\zeta}}})$ we have
$\|D\ell({\bm{y}}^{\prime},{\bm{\zeta}}^{\prime})\|\leq K_{{\bm{y}},{\bm{\zeta}}}\quad\mbox{ and }\quad\|D^{2}\ell({\bm{y}}^{\prime},{\bm{\zeta}}^{\prime})\|\leq K_{{\bm{y}},{\bm{\zeta}}}\,,$
where $D\ell$ and $D^{2}\ell$ denote the first and second derivative in the second argument. So for all such ${\bm{y}}^{\prime}\in B({\bm{y}},1/K_{{\bm{y}},{\bm{\zeta}}})$ and ${\bm{\zeta}}^{\prime},{\bm{\zeta}}^{\prime\prime}\in B({\bm{\zeta}},1/K_{{\bm{y}},{\bm{\zeta}}})$ we have
$|\ell({\bm{y}}^{\prime},{\bm{\zeta}}^{\prime})-\ell({\bm{y}}^{\prime},{\bm{\zeta}}^{\prime\prime})|\leq K_{{\bm{y}},{\bm{\zeta}}}\|{\bm{\zeta}}^{\prime}-{\bm{\zeta}}^{\prime\prime}\|\quad\mbox{ and }\quad|D\ell({\bm{y}}^{\prime},{\bm{\zeta}}^{\prime})-D\ell({\bm{y}}^{\prime},{\bm{\zeta}}^{\prime\prime})|\leq K_{{\bm{y}},{\bm{\zeta}}}\|{\bm{\zeta}}^{\prime}-{\bm{\zeta}}^{\prime\prime}\|\,.$
Cover the set $\{({\bm{y}},{\bm{\zeta}}):\|{\bm{y}}\|\leq C,\|{\bm{\zeta}}\|\leq R\}$ with the balls $\cup_{{\bm{y}}}B({\bm{y}},1/K_{{\bm{y}},{\bm{\zeta}}})$. By compactness, there is a finite subcover $({\bm{y}}_{1},{\bm{\zeta}}_{1}),\ldots,({\bm{y}}_{r},{\bm{\zeta}}_{r})$, so we can take $K_{R}=\max_{i\in[r]}K_{{\bm{y}}_{i},{\bm{\zeta}}_{i}}<\infty$ and the lemma holds since $\|{\bm{y}}\|\leq C$ almost surely by Assumption 2.1. ∎
Lemma D.6.
The function $h({\bm{x}};\cdot)$ is locally bounded, Lipschitz and smooth in its second argument: for any $R>0$ there exists $K_{R}$ such that for any ${\bm{\psi}},{\bm{\psi}}^{\prime}\in B(0,R)$,
$\|h({\bm{x}};{\bm{\psi}})\|$ $\leq K_{R}$
$\|h({\bm{x}};{\bm{\psi}})-h({\bm{x}};{\bm{\psi}}^{\prime})\|$ $\leq K_{R}\|{\bm{\psi}}-{\bm{\psi}}^{\prime}\|$
$\|Dh({\bm{x}};{\bm{\psi}})-Dh({\bm{x}};{\bm{\psi}}^{\prime})\|$ $\leq K_{R}\|{\bm{\psi}}-{\bm{\psi}}^{\prime}\|\,,$
almost surely over ${\bm{x}}$. Here $Dh({\bm{x}},\cdot)\in\mathbb{R}^{d_{out}}\times R^{p}$ is the derivative in the second argument.
Proof.
Analogous to proof of Lemma D.5, using continuous twice-differentiability of $h$ and boundedness of $\|{\bm{x}}\|$. ∎
Lemma D.7 (Local Lipschitzness of loss and loss derivative).
When ${\bm{\theta}}=({\bm{u}},{\bm{v}})\in\mathbb{R}^{2p}$ and $f_{\mathsf{NN}}({\bm{x}};{\bm{\theta}})=h({\bm{x}};{\bm{u}}\odot{\bm{u}})$ the following holds for ${\bm{g}}({\bm{\theta}})$ defined in (4). For any $R>0$, there exists $K_{R}<\infty$ such that for any ${\bm{\theta}},{\bm{\theta}}^{\prime}\in B(0,K_{R})$,
$\|{\bm{g}}({\bm{\theta}})-{\bm{g}}({\bm{\theta}}^{\prime})\|$ $\leq K_{R}\|{\bm{\theta}}-{\bm{\theta}}^{\prime}\|\,$
$\|\nabla_{{\bm{\theta}}}\mathcal{L}({\bm{\theta}})-\nabla_{R}\mathcal{L}({\bm{\theta}}^{\prime})\|$ $\leq K_{{\bm{\theta}}}\|{\bm{\theta}}-{\bm{\theta}}^{\prime}\|$
$|\mathcal{L}({\bm{\theta}})-\mathcal{L}({\bm{\theta}}^{\prime})|$ $\leq K_{R}\|{\bm{\theta}}-{\bm{\theta}}^{\prime}\|\,.$
Proof.
Let ${\bm{\theta}}=({\bm{u}},{\bm{v}}),{\bm{\theta}}^{\prime}=({\bm{u}}^{\prime},{\bm{v}}^{\prime})$. This follows immediately from the local Lipschitzness and smoothness of $h$ and $\ell$ in Lemmas D.5 and D.6, as well as
$\|{\bm{g}}({\bm{\theta}})-{\bm{g}}({\bm{\theta}}^{\prime})\|$ $=\|\operatorname{\mathbb{E}}_{{\bm{x}},{\bm{y}}}[Dh({\bm{x}};{\bm{u}}\odot{\bm{v}})^{\top}D\ell({\bm{y}},h({\bm{x}};{\bm{u}}\odot{\bm{v}}))^{\top}-Dh({\bm{x}};{\bm{u}}^{\prime}\odot{\bm{v}}^{\prime})^{\top}D\ell({\bm{y}},h({\bm{x}};{\bm{u}}^{\prime}\odot{\bm{v}}^{\prime}))^{\top}]\|\,.$